### v1

# customstyle linecolor=gray!100, linewidth=3pt, innerleftmargin=3pt, topline=false, rightline=false, bottomline=false, leftline=true, innerrightmargin=3pt, innertopmargin=3pt, innerbottommargin=3pt, backgroundcolor=gray!15

Abstract.
Atomic commits, each of which addresses a single development concern, are a best practice in software development. However, developers frequently produce tangled commits that mix unrelated changes due to practical constraints or unclear boundaries, negatively impacting code review and maintenance. Although prior commit untangling approaches—rule-based, feature-based, or graph-based—have made progress, they often rely on shallow signals and fail to distinguish between explicit dependencies (e.g., control/data flow) and implicit ones (e.g., semantic or conceptual relationships). In this paper, we propose ColaUntangle, a new collaborative consultation framework for commit untangling that models both explicit and implicit dependencies among code changes. ColaUntangle integrates Large Language Model (LLM)-driven agents in a multi-agent architecture: one agent specializes in explicit dependencies, another in implicit ones, and a reviewer agent synthesizes their perspectives through iterative consultation. To capture explicit and implicit contextual information, we construct multi-version Program Dependency Graphs ($\delta$-PDG), enabling agents to reason over code relationships with both symbolic and semantic depth. We evaluate ColaUntangle on two widely-used datasets (1,612 C# and 14k Java tangled commits). Experimental results show that ColaUntangle outperforms the best-performing baseline, achieving an improvement of 44% on the C# dataset and 100% on the Java dataset. These findings highlight the potential of LLM-based collaborative frameworks for advancing automated commit untangling tasks.
Commit untangling, Atomic commits, Explicit and implicit dependencies, LLMs, Multi-agent collaboration
1. Introduction
In collaborative software development, a commit is the basic unit of code change, consisting of source file modifications and a message summarizing their purpose (shen2021smartcommit, ). Ideally, each commit should be atomic, addressing a single concern—such as implementing a feature, fixing a bug, or refactoring (li2022utango, ; guo2019decomposing, ). Atomic commits are a widely endorsed best practice for enhancing readability, facilitating reviews, and aiding maintenance by clearly linking changes to their intent (tao2012empiricalwork, ; rigby2012contemporary, ; wang2019cora, ; herbold2022fine, ). As a result, many software communities and companies recommend small, single-purpose commits (fan2024HD-GNN, ; linuxSubmittingPatches, ; microsoftGitGuidance, ).
However, in practice, developers often create tangled commits, which combine code changes related to multiple development concerns within a single commit (fan2024HD-GNN, ; murphy2011we, ). Tangled commits commonly arise due to time constraints or unclear boundaries between concerns during implementation (fan2024HD-GNN, ; parțachi2020flexeme, ; li2022utango, ). For example, a developer may simultaneously refactor code while fixing a bug alongside documentation updates, resulting in a commit that does not adhere to the principle of separation of concerns (herzig2016impact, ; herzig2013impact, ; nguyen2013filtering, ).
Empirical studies (herzig2013impact, ; herzig2016impact, ; tao2015partitioning, ; nguyen2013filtering, ) find that tangled commits are prevalent, indicating that 11% to 40% of commits in software repositories are tangled. Such commits hinder program comprehension, complicate maintenance (tao2012empiricalwork, ; parțachi2020flexeme, ; li2015gated, ). Moreover, tangled commits reduce the accuracy of automated tools such as bug prediction and localization models that rely on mining repository history (herzig2013impact, ; herzig2016impact, ; zhou2019devign, ; kirinuki2014hey, ). These tools often assume that each commit addresses a single concern, so tangled commits introduce noise and degrade model performance.
Researchers have proposed various approaches to untangling commits to split a tangled commit into separate coherent changes. These methods fall into three categories based on their automation level: heuristic rule-based, feature-based, and graph clustering-based. Heuristic methods use manually defined rules—such as line distance, textual similarity, or co-change frequency—to assess change relatedness (barnett2015Heuristic-rule, ; herzig2013impact, ; kirinuki2016splitting, ; wang2019cora, ). Feature-based methods employ handcrafted features (e.g., same method, class, or package) in supervised models to classify change pairs (dias2015EpiceaUntangler, ; yamashita2020changebeadsthreader, ). Graph clustering-based methods build program graphs (e.g., PDGs) to model structural dependencies and use clustering or node embeddings to group related changes (parțachi2020flexeme, ; li2022utango, ; xu2025COMUNT, ; fan2024HD-GNN, ).
Despite their methodological differences, these approaches share several key limitations. First, they often rely on surface-level signals or structural proximity and lack the capacity for deeper semantic reasoning. Graph-based methods attempt to encode some semantic relationships via graph embeddings, but they typically require substantial training and are limited in their ability to generalize across projects or languages. Second, most approaches act as black-box models, offering limited interpretability and little insight into why specific code changes are grouped together. Third and most critically, prior work does not clearly distinguish or integrate different types of dependencies among code changes. For instance, two edits may be connected via control flow (explicit) or may reflect a coherent conceptual refactoring (implicit), yet existing models either fail to recognize such links or treat them uniformly, resulting in reduced effectiveness on complex or ambiguous commits.
We argue that effective commit untangling must go beyond surface-level patterns and rely on a principled framework that explicitly accounts for both explicit and implicit dependencies. Explicit dependencies refer to observable relationships such as control or data flow, containment, or static code references. Implicit dependencies involve semantic connections—such as conceptual similarity, logical association, or shared intent—that are not necessarily reflected in the program structure. While some existing methods may partially capture these relationships through learned embeddings or co-change patterns, they do not explicitly model or directly leverage these dependencies as integral elements of the untangling process.
Recent advances in large language models (LLMs) have demonstrated strong performance in software engineering tasks that require semantic understanding, reasoning, and explanation (yang2024harnessing, ; fan2023large, ; hou2024large, ). LLMs are particularly well-suited to identify implicit dependencies that go beyond what is detectable via static analysis or statistical patterns. However, a single LLM acting alone may not consistently balance different dependency perspectives or resolve ambiguous cases. Inspired by recent work on multi-agent collaboration (he2025llm, ; wang2025colacare, ; tang2023medagents, ; Kim2024mdagents, ), we propose decomposing the untangling task into subtasks, each handled by specialized agents, and enabling them to collaborate through consultation.
To this end, we introduce ColaUntangle, a collaborative consultation framework for commit untangling that integrates both explicit and implicit dependencies. We construct multi-version Program Dependency Graphs ($\delta$-PDG) to capture structural and contextual information for code changes. Then, we design a multi-agent architecture driven by LLMs: an explicit worker agent focused on explicit dependencies (e.g., control/data flow), an implicit worker agent focused on implicit dependencies (e.g., conceptual relationships), and a reviewer agent that synthesizes and reconciles their results. All agents provide not only untangling decisions but also explanations. Through iterative interaction, these agents simulate human-like consultation and collectively decide on the final untangling outcome.
To evaluate the effectiveness and efficiency of the proposed approach, we conduct experiments on the widely-used C# and Java datasets (parțachi2020flexeme, ; li2022utango, ; fan2024HD-GNN, ) with 1,612 and 14k tangled commits. The evaluation results show that ColaUntangle achieves better untangling results (i.e., 44% and 100% enhancement of effectiveness for C# and Java, compared to the best-performing baseline). Overall, our work makes the following contributions:
  • ColaUntangle: a collaborative consultation model leveraging LLM-driven agents for commit untangling. We propose ColaUntangle, the first LLM-driven collaborative consultation model that untangles commits by considering explicit and implicit dependencies among code changes, simulating human collaborative decision-making to generate results and explanations.
  • Definition of explicit and implicit dependencies for commit untangling. We define a clear and generalizable set of explicit and implicit dependencies between code changes, providing principled guidance that enables both automated and explainable commit untangling.
  • Extensive empirical evaluation. We evaluate ColaUntangle against the state-of-the-art approaches, and the results demonstrate its superior performance. Our tool and data are available at https://anonymous.4open.science/r/ColaUntangle-0109.
2. Motivation and Key Ideas
2.1. Motivating Examples
The following real-world examples serve to motivate ColaUntangle and to introduce its fundamental concepts.
Figure 1. Motivating Examples
Example 1. Figure 1A illustrates the changes made in commit 1ad6dbc of the Elasticsearch project. This atom commit updates the index used in the testMultipleBatchesWithLookupJoin function from "idx" to "no_sort_field_idx". The commit includes three separate code changes: lines 3–4 create the new index, lines 6–9 query using the updated index, and lines 11–12 delete the index. These changes do not exhibit any obvious data or control dependencies, nor do they share structural similarities in their syntax. However, they are semantically related, as they collectively form a complete and independent development activity involving the creation, usage, and deletion of an index. This semantic coherence indicates that they belong to a single concern. Existing untangling approaches may fail to handle such commits because they primarily rely on explicit data dependencies or structural similarities between code changes, overlooking the semantic relationships that often define a developer’s intent.
Observation 1 [Semantic]. The semantics of code changes are crucial for determining their underlying concern. Changes that belong to the same concern often share semantic similarity.
Example 2. Figure 1B shows another atom commit. Lines 3–6 define an enum class representing ordering options, and lines 8–16 implement a sorting method that utilizes this enum (only the if block is shown in the figure, as the else block has a similar structure). The code changes in lines 3–6 and lines 8–16 exhibit logical dependencies respectively, as they belong to the same class or method implementation. Furthermore, there is a clear data dependency: line 11 directly uses the value defined in the enum class introduced in lines 3–6. Due to these logical and data dependencies, these changes are categorized as belonging to a single concern.
Observation 2 [Data and control dependency]. Code changes within the same concern often exhibit data or control dependencies.
Example 3. Figure 1C shows a tangled commit containing two distinct concerns: ”remove LicensingCharset” and ”reformat code”. The first concern (at lines 3 and 6) involves changes to the charset used in the TrialLicenseUtils class (only a subset of these changes is shown, as other charset modifications exhibit similar structure). Each of these changes is identical, replacing LicensesCharset with StandardCharsets. The affected statements are cloned code segments used to support different variables (issueTo and feature), resulting in high similarity in both textual content and code structure.
Observation 3 [Code similarity]. Code changes within the same concern often exhibit high similarity in text and structure.
The second concern focuses on reformatting code at lines 9, and 13, spanning two files. These changes represent typical cosmetic edits, which include syntactic formatting, refactoring, or non-functional textual modifications. Such edits are common in daily software development and usually belong to a single concern. However, existing approaches fail to detect and classify these changes correctly, as their rules, features, or graph representations do not model cosmetic edits.
Observation 4 [Cosmetic edit]. Code changes caused by cosmetic edits, such as syntactic formatting, refactoring, or non-functional textual modifications, often belong to the same concern.
2.2. Key Ideas
Inspired by the above observations, we propose ColaUntangle with the following key ideas.
2.2.1. Key Idea 1 [Explicit and Implicit Dependency]
Based on Observations 1–4, we find that code changes can be grouped into different concerns due to various factors. To provide comprehensive guidance for our method, we categorize the characteristics of code changes belonging to the same concern into two types: explicit dependencies and implicit dependencies.
Explicit dependencies capture the direct relationships between code changes, such as data or control dependencies. In contrast, implicit dependencies capture the underlying reasons why code changes without obvious structural or data relationships may still belong to the same concern, such as semantic similarity or shared development purpose. Specifically, we define the following rules:
Explicit Dependencies Rules:
  • Code changes with data dependencies often belong to the same concern. [Observation 2]
  • Code changes with control dependencies often belong to the same concern. [Observation 2]
Implicit Dependencies Rules:
  • Code changes with semantic similarity may belong to the same concern. [Observation 1]
  • Code changes with high textual or structure similarity may belong to the same concern. [Observation 3]
  • Code changes introduced for cosmetic edits, such as syntactic formatting, refactoring, or non-functional textual modifications, often belong to the same concern. [Observation 4]
2.2.2. Key Idea 2 [LLM-driven Foundation Models]
Building on Observation 1, 3 and 4, we decide to leverage the powerful capabilities of LLMs rather than traditional machine learning or neural network models to untangle commits. LLMs exhibit superior performance in comprehending code semantics, performing reasoning, and adapting to diverse decision-making tasks (liu2024large, ; jin2024llms, ). By employing LLM-driven agents, ColaUntangle can effectively understand the intent and implicit dependencies underlying code changes beyond purely structural or syntactic similarities. This enables our approach to untangle commits with a deeper semantic understanding which is difficult to achieve based on heuristic rules, feature representations and graph models.
2.2.3. Key Idea 3 [Collaborative Consultation]
As seen in Observations 1–4, code changes may belong to different concerns due to diverse and sometimes conflicting factors, making it difficult to define a single generalized untangling criterion. Therefore, we design a multi-agent collaborative consultation process comprising distinct roles and iterative negotiation. Specifically, we define two types of worker agents: one specializing in analyzing explicit dependencies between code changes and the other focusing on implicit dependencies. Additionally, we introduce a reviewer agent that synthesizes and integrates the outputs from both worker agents. These three agents engage in multiple rounds of collaborative consultation, producing the final untangling result upon reaching consensus.
3. Methodology
We propose ColaUntangle, a collaborative framework that leverages LLM-driven agents to untangle commits by identifying both explicit and implicit dependencies among code changes. As shown in Figure 2, ColaUntangle first extracts structured code change information, including commit diffs and contexts derived from $\delta$-PDG. It then employs a multi-agent consultation process, where agents iteratively analyze dependencies and synthesize untangling decisions until reaching consensus.
Figure 2. Overall Workflow of ColaUntangle
Figure 3. Code Change Example with Explicit and Implicit Contexts Derived from $\delta$-PDG
3.1. Structured Code Change Information
To enrich the contextual information available for dependency analysis, we first extract explicit contexts and implicit contexts from the multi-version program dependency graph ($\delta$-PDG).
A program dependency graph (PDG) (ferrante1987PDG, ) is a directed graph that consists of a set of nodes N and a set of edges E, where each node n $\in$ N represents a program statement or a conditional expression; and each edge e $\in$ E represents the data flow or control flow among the statements. However, PDG can only describe the dependency relationships between statements of individual program. In order to simultaneously reflect the statement dependencies before and after program modification, we also need to introduce $\delta$-PDG. The multi-version program dependency graph ($\delta$-$PDG^{i,j}$) (parțachi2020flexeme, ) is the disjoint union of all nodes and edges across all versions in the PDGs at versions i and j.
For most modification operations, the changed parts account for only a small proportion of the entire program. To reduce unnecessary information, we introduce the concept of explicit contexts. The explicit contexts is a compressed $\delta$-PDG graph obtained by retaining only the modified nodes and their relationships in the $\delta$-PDG. Although the explicit contexts loses some context information, it still retains the data dependencies and control dependencies between the changed nodes. Consequently, the implicit contexts retains only the modified nodes in the $\delta$-PDG and their neighbor nodes in one hop. Compared with the explicit contexts, the implicit contexts retains more surrounding code information.
Figure 3 illustrates a code change scenario involving two statement replacements (at lines 2 and 6). Figure 3 (a) shows the $\delta$-PDG of this code snippet, where red nodes represent the statements before the change, green nodes represent the modified statements, and gray nodes denote the unchanged statements. Figure 3 (b) presents the explicit contexts, which are obtained by compressing the $\delta$-PDG in Figure 3 (a). Specifically, we retain the changed nodes and their direct dependency edges; additionally, if two changed nodes lack a direct dependency but are connected via a dependency path (channel), we add a compressed edge between them. For example, nodes 2 and 6 have no direct dependency but are connected via a path through nodes 4 and 5, resulting in a compressed edge between them. This compressed edge is labeled with the variables involved in the data flow along the original dependency path. Figure 3 (c) shows the implicit contexts, which retain both the changed nodes and their immediate neighbors (within one-hop dependencies). In this example, node 9 is excluded because it does not have a direct dependency with any of the changed nodes.
3.2. Multi-Agent Collaborative Consultation
As shown in Figure 2, the multi-agent collaborative consultation involves three types of agents: two worker agents specializing in mining explicit and implicit dependencies between code changes, respectively, and a reviewer agent responsible for reviewing and synthesizing the untangling results. The process begins with the two worker agents generating an initial untangling of a tangled commit based on its structured code information. The reviewer agent then synthesizes these results, initiating the collaborative consultation. In each iteration, the worker agents provide feedback on the current synthesized result, which the reviewer agent uses to revise synthesized result accordingly. This iterative process continues until the agents reach a consensus or a maximum number of rounds is reached.
3.2.1. Agent Building
Figure 4. Worker Agent
Figure 5. Reviewer Agent
ColaUntangle builds three types of agents. The first two are worker agents, each specializing in untangling commits based on either explicit or implicit dependencies between code changes, as illustrated in Figure 4. The third is a reviewer agent, which specializes in reviewing and synthesizing the results from the worker agents, as shown in Figure 5.
Worker Agent. Based on the explicit and implicit dependency rules summarized in Section 2.2, we define two types of worker agents, as illustrated in Figure 4. The worker agents specialize in untangling commits by analyzing explicit and implicit dependencies between code changes, respectively. We refer to the explicit worker agent as EA and the implicit worker agent as IA. We define the roles of these agents using the following role definition prompt: “You are an expert in untangling tangled commits, specializing in analyzing explicit (or implicit) dependencies between code changes and comprehending their underlying semantic relationships”. We also provide each agent with the corresponding explicit or implicit dependency rules (concluded in Section 2.2.1) to guide their analysis. Each worker agent has two main functions. The first function is “untangle”. EA untangles tangled commits based on the inputs of commit diffs and explicit contexts, while IA performs untangling based on commit diffs and implicit contexts. Both agents are required to generate explanations for their decisions to facilitate understanding of their underlying criteria. The second function is “validate”. In this function, each worker agent compares the synthesized result generated by the reviewer agent (described in the next section) with its own untangling result and provides its opinion on the synthesized result. This includes indicating agreement or disagreement and, if disagreeing, providing a revised untangling result along with an explanation.
Reviewer Agent. The untangling results produced by the two worker agents may conflict due to their different expertise and the presence of diverse explicit and implicit dependencies within code changes. To resolve these conflicts, we define a reviewer agent, referred to as RA, which specializes in reviewing and synthesizing the untangling results from the worker agents, as illustrated in Figure 5. We define the reviewer agent’s role using the following role definition prompt: “You are an expert reviewer specializing in reviewing and synthesizing untangling results from worker agents. You have deep expertise in analyzing both explicit and implicit dependencies between code changes”. To support its role, we provide RA with both explicit and implicit dependency rules to guide its synthesis process. RA has three main functions: synthesize, revise and stop. In the synthesize function, RA synthesizes the results from EA and IA based on the structured code change information (commit diffs, explicit contexts, and implicit contexts). In the revise function, RA revises its synthesized result based on the new opinions generated by the two worker agents during validation. Finally, in the stop function, when the collaborative consultation process reaches its maximum iteration limit, RA produces the final untangling result. For each function, RA is required to provide clear explanations for its decisions.
Algorithm 1 Collaborative Consultation
1:Experts $EA$, $IA$, $RA$; initial synthesized result $r_{RA}$; maximum attempts $t$; prompt template $Prompt$
2:Final result $r_{f}$
3:// Initialize variables
4:$consensus\leftarrow False$, $round\leftarrow 0$
5:// Iterative consultation
6:while $consensus=False$ and $round<t$ do
7:    $round\leftarrow round+1$
8:    $consensus\leftarrow True$
9:    $EA\_opinion\leftarrow EA.validate(Prompt(r_{RA}))$
10:    $IA\_opinion\leftarrow IA.validate(Prompt(r_{RA}))$
11:    if $EA\_opinion.agree=False$ or $IA\_opinion.agree=False$ then
12:        $consensus\leftarrow False$
13:        if $round=t$ then
14:           $r_{RA}\leftarrow RA.stop(EA\_opinion,IA\_opinion)$
15:        else
16:           $r_{RA}\leftarrow RA.revise(EA\_opinion,IA\_opinion)$
17:        end if
18:    end if
19:end while
20:$r_{f}\leftarrow r_{RA}$
21:return $r_{f}$
3.2.2. Utangling Workflow
The workflow of ColaUntangle is illustrated in Figure 2. The collaboration begins with the worker agents (EA and IA) providing their initial untangling results. Subsequently, the reviewer agent (RA) synthesizes these results into a unified untangling outcome. This initiates an iterative collaborative consultation process, during which EA and IA express their opinions on the current synthesized result. RA then considers their feedback and revises the synthesized result accordingly. This iterative process continues until EA and IA reach a consensus on RA’s synthesized result or a predefined maximum number of rounds is reached. Specifically, the untangling workflow comprises three phases: generation of initial reviews, synthesis of preliminary untangling results, and collaborative consultation process.
In the first phase, generation of initial reviews, EA and IA are instructed to generate their preliminary untangling results $r_{EA}$ and $r_{IA}$ using their Untangle function, based on the commit diffs ($Info_{diffs}$) and the explicit ($Info_{EC}$) or implicit ($Info_{IC}$) contexts:
$r_{EA}=EA.Untangle(Prompt(Info_{diffs},Info_{EC})$
$r_{IA}=IA.Untangle(Prompt(Info_{diffs},Info_{IC})$
In the second phase, synthesis of preliminary untangling results, RA uses its Synthesize function to review and integrate the results from EA and IA into a synthesized result $r_{RA}$
$r_{RA}=RA.Synthesize(Prompt(r_{EA},r_{IA},Info_{diffs},Info_{EC},Info_{IC})$
The final phase, collaborative consultation process, is iterative as shown in Algorithm 1. During each round $j$, the worker agents EA and IA evaluate the current synthesized result and express their agreement or disagreement. In cases of disagreement, they are required to provide detailed rationales. To be specific, in round $j$-th:
$EA\_opinion^{j}=EA.validate(Prompt(r_{RA}^{j-1}))$
$IA\_opinion^{j}=IA.validate(Prompt(r_{RA}^{j-1}))$
If both agents reach unanimous agreement, the iterative process terminates (next loop condition is not satisfied). Otherwise, if the maximum round limit has not been reached, RA uses its Revise function to consider the opinions from EA and IA, refining the current synthesized result to produce a new synthesized result accordingly:
$r_{RA}^{j}=RA.revise(Prompt(EA\_opinion^{j},EA\_opinion^{j})$
If the maximum round limit is reached without achieving consensus, RA uses its Stop function to finalize the synthesized result based on the current results and the agents’ opinions:
$r_{RA}^{j}=RA.stop(Prompt(EA\_opinion^{j},EA\_opinion^{j})$
In the next round, the worker agents once again express their opinions towards the new synthesized result and try to convince the RA. The RA considers their statements and further revises the result until all worker agents reach a consensus or the maximum number of rounds is reached.
4. Experimental Setups
4.1. Datasets
We use two datasets that have been widely used in previous work: 1) A C# dataset used in (chen2022attributedgraph, ; herzig2016impact, ; parțachi2020flexeme, ; li2022utango, ; fan2024HD-GNN, ), which contains 1,612 tangled commits (each of which has $\leq$ 3 concerns) across 9 GitHub projects. 2) A Java dataset used in (li2022utango, ; fan2024HD-GNN, ). The Java dataset contains over 14k tangled commits in 10 GitHub repositories, where the number of concerns in a commit ranges from 2-32. Both datasets were constructed by first selecting atomic commits based on criteria such as temporal proximity, namespace similarity, and file coupling, and then artificially tangling them via git cherry-picking to simulate realistic tangled commit scenarios.
4.2. Baselines
Commit Untangling-specfic Baselines. The Baselines used for C# dataset are:
  • Barnett et al. (barnett2015Heuristic-rule, ): A heuristic rule-based approach considering def-use, use-use, and same-enclosing-method relations among code changes.
  • Herzig et al. (herzig2016impact, ): A heuristic rule-based approach combine various Confidence Voters and build triangle partition matrix to untangle the commits.
  • Flexeme (parțachi2020flexeme, ): A graph clustering-based approach builds a $\delta$-NFG (proposed in Flexeme) of commits and then applies agglomerative clustering on this $\delta$-NFG to untangle the commits.
  • $\delta$-PDG+CV (parțachi2020flexeme, ): A variant of Flexeme by applying Herzig et al.’s confidence voters directly to $\delta$-PDG (proposed in Flexeme).
  • UTango (li2022utango, ): A Graph clustering-based approach builds a $\delta$-PDG from commits and then applies GNN and agglomerative clustering on this $\delta$-PDG to untangle the commits.
  • HD-GNN (fan2024HD-GNN, ): A Graph clustering-based approach builds a entity reference graph from commits and then applies GNN to aggregate the embeddings of nodes and edges in the graph and classify the relation between the code changes.
The baselines used for Java dataset are: Barnett et al., Herzig et al., UTango, and HD-GNN mentioned above, as all these approaches support analysis of Java program files. SmartCommit is a graph clustering-based interactive method that builds a Diff Hunk Graph and applies a graph partitioning algorithm to untangle commits. We also include SmartCommit and its baselines, Base-1 (a rule-based approach that groups all changes into a single concern) and Base-2 (which groups changes by file), in our comparison.
LLM-based Baselines. We incorporate baselines utilizing single LLM for both C# and Java datasets. We employ two different prompt strategies: zero-shot and zero-shot CoT. Zero-shot refers to prompting LLM to perform a task without providing examples, zero-shot CoT directly incorporates the prompt Let’s think step by step after a question to facilitate inference (kojima2022zeroshot, ).
4.3. Metrics
We use two evaluation metrics: $Accuracy^{c}$ proposed by Utango (li2022utango, ) and $Accuracy^{a}$ proposed by Flexeme (parțachi2020flexeme, ). $Accuracy^{c}$ is defined as the percentage of the changed statements that are labeled with a correct cluster/concern in all the statements in a commit:
$Accuracy^{c}=\frac{\#changed\ stmts\ w.\ correct\ clusters}{\#all\ changed\ stmts\ in\ commit}$
$Accuracy^{a}$ is similar to $Accuracy^{c}$ except that it considers all statements:
$Accuracy^{a}=\frac{\#stmts\ w.\ correct\ clusters}{\#all\ stmts\ in\ commit}$
4.4. Implementation Details
We utilize the widely adopted and publicly available DeepSeek-V3 model (liu2024deepseek, ) from the Azure OpenAI service [^1] https://learn.microsoft.com/en-us/azure/ai-services/openai/ as the reasoning engine in our experiments, which is also used for the LLM-based baselines. All experiments are conducted under a zero-shot setting. From an initial evaluation on 50 randomly selected tangled commits, we observed an average of 1.27 consultation rounds to reach consensus. Thus, we set the maximum number of rounds $t$ to 3, aligning with ColaCare (wang2025colacare, ) (a medical multi-agent collaborative consultation model), and balancing iterative refinement with computational efficiency. In terms of computational cost, our approach incurs approximately \$2.92 per 100 examples (around ¢2.92 per example), with an average inference time of 125.13 seconds per example [^2] We acknowledge that our approach requires longer inference time compared to prior methods. However, it eliminates the need for manual heuristic rule design, feature engineering, and model training, making it relatively cost-effective given its significant improvement in untangling accuracy..
5. Experimental Results
To evaluate ColaUntangle, we seek to answer the following research questions:
  • [RQ1. Overall Performance] How well does ColaUntangle perform in comparison with the state-of-the-art commit untangling approaches?
  • [RQ2. Ablation Study] What is the contribution of each proposed module to the performance?
  • [RQ3. Sensitivity to Different LLMs] How does ColaUntangle’s performance vary with different LLMs?
Table 1. Performance on the C# Dataset ($Accuracy^{c}$%/$Accuracy^{a}$%)
Method CL CM HF HU LE NA NJ HI RS Overall
Previous Untangling Approaches
Barnett et al. 14/20 */20 12/15 13/25 7/16 */9 7/13 10/14 10/13 7/13
Herzig et al. 28/58 */65 28/62 27/53 27/66 */63 28/64 26/57 31/70 28/66
$\delta$-PDG + CV 34/81 */90 36/86 30/69 35/78 */83 34/78 37/94 33/64 35/84
Flexeme 34/87 */70 33/77 33/70 35/80 */87 27/62 32/80 34/86 34/83
UTango 46/90 */* 46/88 44/89 46/89 */* 41/83 46/91 46/88 46/90
HD-GNN 52/91 */* 52/90 50/90 50/91 */* 45/86 54/93 47/90 50/90
ColaUntangle Variants
$LLM_{ZeroShot}$ 62/90 68/90 64/93 44/66 58/92 52/88 68/97 63/91 60/92 60/90
$LLM_{ZeroShotCoT}$ 63/90 68/90 64/93 45/66 59/92 54/88 68/97 65/92 60/92 61/90
ColaUntangleno_comments 74/94 74/92 80/96 65/76 70/94 58/90 66/97 68/93 60/93 70/93
ColaUntangle 75/94 72/90 80/96 65/77 70/94 65/91 75/98 74/93 64/93 72/93
  • * refers no available data point.
    ColaUntangleno_comments refers to run our method on the dataset excluding comments.
Table 2. Performance on the Java Dataset ($Accuracy^{c}$%)
Method SB ES RJ GU RE DU GH ZX DR EB Overall
Previous Untangling Approaches
Barnett et al. 21 27 24 25 19 18 24 20 18 21 17
Herzig et al. 25 32 27 27 25 20 27 26 29 24 23
SmartCommit 29 34 31 33 29 24 33 34 32 27 30
Base-1 14 19 18 21 22 15 22 21 17 13 17
Base-2 23 22 28 31 24 19 27 29 29 22 24
UTango 32 36 33 37 34 30 35 38 34 31 34
HD-GNN 37 40 38 42 36 32 40 42 39 36 38
ColaUntangle Variants
$LLM_{ZeroShot}$ 66 68 66 50 54 60 51 73 66 57 63
$LLM_{ZeroShotCoT}$ 66 70 68 68 62 62 49 70 65 57 66
ColaUntangleno_comments 70 80 69 77 60 75 67 71 79 71 76
ColaUntangle 70 79 70 78 61 75 66 73 80 72 76
  • Since previous studies do not report $Accuracy^{a}$on the Java dataset, we follow their practice and report only $Accuracy^{c}$to ensure a fair comparison.
    ColaUntangleno_comments refers to run ColaUntangle on the dataset excluding comments.
5.1. RQ1: Overall Performance
Table 1 compares metrics on the C# dataset (including nine GitHub projects: CL: Commandline, CM: CommonMark, HF: Hangfire, HU: Humanizer, LE: Lean, NA: Nancy, NJ: Newtonsoft.Json, NI: Ninject, RS: RestSharp). Compared to previous untangling approaches, ColaUntangle outperforms SOTA, improving Barnett et al., Herzig et al., $\delta$-PDG+CV, Flexeme, Utango and HD-GNN by 928.57%, 157.14%, 105.71%, 111.76%, 56.52% and 44.00% for $Accuracy^{c}$; 615.38%, 40.91%, 10.71%, 12.05%, 3.33% and 3.33% for $Accuracy^{a}$. For LLM-Based baselines, ColaUntangle improves $LLM_{ZeroShot}$ and $LLM_{ZeroShotCoT}$ by 16.67% and 16.39% on $Accuracy^{c}$, and by 3.33% on $Accuracy^{a}$ for both. Compared to ColaUntangleno_comments, which excludes comment statements, ColaUntangle achieves a 2.86% higher $Accuracy^{c}$, while $Accuracy^{a}$ remains the same.
Table 2 shows the Java dataset comparison (including ten GitHub projects: SB: spring-boot, ES: elasticsearch, RJ: RxJava, GU: guava, RE: retrofit, DU: dubbo, GH: ghidra, ZX: zxing, DR: druid, EB: EventBus), where ColaUntangle improves Barnett et al., Herzig et al., Base-1, Base-2, SmartCommit Utango and HD-GNN by 347.06%, 230.43%, 153.33%, 347.06%, 216.67%, 123.53% and 100.00% for $Accuracy^{c}$. For LLM-based baselines, ColaUntangle improves over $LLM_{ZeroShot}$ and $LLM_{ZeroShotCoT}$ by 16.67% and 16.39% respectively, and is the same as ColaUntangleno_comments for $Accuracy^{c}$.
The results show that ColaUntangle consistently outperforms all baseline models, with notable improvements in the $Accuracy^{c}$ metric. Importantly, unlike most existing methods, ColaUntangle can untangle comment statements, which are common cosmetic edits in real-world commits, making it more practical for software development. ColaUntangleno_comments also surpasses all prior approaches, indicating that while comments can slightly enhance performance, our framework remains effective even without them.
Regarding collaborative consultation rounds, the average number of rounds is 1.38: 75.50% of cases converge within one round, 11.53% within two rounds, and 13.16% reach the maximum of three rounds. Notably, in 48.91% of evaluated commits, the explicit and implicit worker agents initially produced non-consensual results, underscoring the importance of collaborative consultation for resolving divergent untangling perspectives.
(a) C# Repositories
(b) Java Repositories
Figure 6. Average $Accuracy^{c}$ and Average Number of Changed Statements in Tangled Commits of Each Repository
We also investigate the factors contributing to accuracy differences across repositories. As shown in Figure 6, for both the C# and Java datasets, the purple bars represent the average $Accuracy^{c}$ in each repository, while the red bars indicate the average number of changed statements of tangled commits in each repository. We observe a general trend where $Accuracy^{c}$ decreases as the number of changed statements increases. This may be because abundant changes in the commit cause LLMs forgetting context or untangling commits more finely than humans expect. These findings suggest future work on hierarchical prompting or human-in-the-loop approaches to better align with human expectations.
{mdframed}
[style=customstyle] Answer to RQ1: ColaUntangle demonstrates improvements over prior baselines on both datasets, improving $Accuracy^{c}$ by 44% on the C# dataset and 100% on the Java dataset over the best existing baseline. The results highlights its potential for practical application in untangling commits.
5.2. RQ2: Ablation Study
We conduct an ablation study to evaluate the contribution of each component within ColaUntangle. As shown in Table 3, we define four ablation settings: (1) w/o Info Tools: Removes explicit and implicit contexts extracted from the $\delta$-PDG. (2) w/o Collaborative Consultation: Uses a single LLM model to mine both explicit and implicit dependencies without the multi-agent consultation process. (3) w/o Explicit Worker Agent (W1) and Collaborative Consultation: Retains only the implicit worker agent to perform untangling. Since there is only a single agent remaining, the collaborative consultation process becomes unnecessary, and the reviewer agent is thus excluded. (4) w/o Implicit Worker Agent (W2) and Collaborative Consultation: The converse of (3), retaining only the explicit worker agent.
From the results, we observe that excluding Collaborative Consultation (setting 2) leads to the most significant performance degradation (17.57%) [^3] For the sake of brevity, all mentions of performance degradation in this section refer specifically to a drop in $Accuracy^{c}$, as it better reflects the model’s performance in real-world tasks compared to $Accuracy^{a}$., highlighting its critical role in achieving optimal untangling results. Removing either the explicit or implicit worker agent (settings 3 and 4) also results in significant performance drops (20.27% and 18.92%, respectively). However, it is important to note that these two settings inherently exclude Collaborative Consultation due to the presence of only a single agent, making their performance degradation partly attributable to the absence of collaborative consultation. When we isolate the effect by excluding the performance drop caused by the lack of collaborative consultation, settings 3 and 4 show additional reductions of only 2.70% and 1.35%, respectively. This indicates that while incorporating both explicit and implicit worker agents enhances performance, the collaborative consultation mechanism itself contributes the most substantial improvement in our framework. Removing Info Tools (setting 1) results in a performance decrease of 9.46%, suggesting that the structured context improves our model’s untangling effectiveness. However, compared to the substantial improvements brought by collaborative consultation, Info Tools mainly serves as foundational support by providing essential contextual information rather than driving the core reasoning and decision-making process.
{mdframed}
[style=customstyle] Answer to RQ2: The ablation results suggest that while structured context information and the integration of both explicit and implicit worker agents supports untangling accuracy, the collaborative consultation mechanism play a more critical role in achieving the effectiveness of ColaUntangle.
Table 3. Ablation Study Results for Each Module ($Accuracy^{c}$%/$Accuracy^{a}$%)
Method C# Dataset Java Dataset Overall
w/o Info Tools 64/91 70/92 67/92
w/o Collaboraive 60/90 63/92 61/91
w/o W1 w/o Collaborative 57/89 63/91 59/90
w/o W2 w/o Collaborative 58/90 63/92 60/91
ColaUntangle 72/93 76/94 74/94
  • W1(2) refer to worker agent that expert in discovering explicit (implicit) dependency between code changes.
5.3. RQ3: Sensitivity to Different LLMs
To evaluate the generalization capability of our framework, we include both closed-source proprietary models (GPT-4o, GPT-4o-mini from OpenAI, Claude-4-sonnet from Anthropic) and open-source frontier models (DeepSeek-V3 and Qwen3-235b) in our comparison. We compare ColaUntangle instantiated with these five models respectively. The performance presented in Table 4 shows that all these LLMs can effectively untangle commits, with DeepSeek-V3, Claude-4-sonnet and GPT-4o-mini performing slightly better. This suggests that while ColaUntangle is generally robust to the underlying LLM choice, stronger reasoning models still yield marginal performance gains. Furthermore, we analyze the average number of collaborative rounds required for consensus: DeepSeek-V3 (1.38), GPT-4o-mini (1.65), GPT-4o (1.35), Qwen3-235b (1.39), and Claude-4-sonnet (1.77). Notably, DeepSeek-V3 not only attains leading accuracy but also requires fewer consensus rounds compared to most competing models (except GPT-4o), highlighting its efficiency in balancing performance and computational cost.
{mdframed}
[style=customstyle] Answer to RQ3: Our evaluation demonstrates that while all tested LLMs effectively handle commit untangling, DeepSeek-V3 achieves the optimal balance of accuracy and efficiency (fewer consensus rounds).
Table 4. Performance of Different LLMs ($Accuracy^{c}$%/$Accuracy^{a}$%)
Method C# Dataset Java Dataset Overall
GPT-4o-mini 72/93 71/93 72/93
GPT-4o 68/92 74/94 71/93
Qwen3-235b 71/93 73/93 72/93
Claude-4-sonnet 70/93 76/94 73/93
ColaUntangle (DeepSeek-V3) 72/93 76/94 74/94
6. Discussion of Error Cases
Figure 7. Examples of Error Cases Related to Fine Grouping Granularity
Figure 8. Examples of Error Cases Related to Coarse Grouping Granularity
In the RQ1 experiments, although ColaUntangle performed well overall, it still misclassified some cases. To support future improvements, we conducted a manual analysis following (fan2024HD-GNN, ), randomly sampling 50 misclassified cases. As shown in Table 5, we grouped the errors into two main types: fine and coarse grouping granularity. Fine grouping granularity refers to ColaUntangle untangles a tangled commit into concerns at finer granularity than the standard answer. In contrast, coarse grouping granularity refers to ColaUntangle untangles a tangled commit into concerns at coarser granularity than the standard answer. Further analysis identified five sub-types that specify the scenarios of these errors.
Table 5. Type of Error Cases
Type Sub-Type Number
Fine Grouping Granularity Judgment Error 14
Composite Commit 17
File Formatting 3
Coarse Grouping Granularity Similar Change 11
Mergeable Logic 5
For fine grouping granularity, we observed three representative sub-types: judgment errors (14 cases), composite commit (17 cases), and file formatting (3 cases). Judgment errors (14 cases) occur when ColaUntangle misinterprets the relationship between changes. For example, in Figure 7 A, changes across two files belong to a single concern related to signature revision, but ColaUntangle misclassified them into two concerns by treating lines 9–10 as an isolated function call modification (configure) unrelated to the other file. Similarly, in Figure 7 B, changes at lines 3–5 and 8–9 belong to the same concern (add more logging to jgroups disco), but ColaUntangle separated them because it focused on different logger methods (debug vs. warn). Composite commit (17 cases) occur when the standard atom commit in the dataset actually contains multiple unrelated activities. For instance, Figure 7 C shows a commit that modifies comments for the create function (lines 3–5) and changes the signature of canAcquire, with the two code changes separated by over 240 lines and lacking any logical connection. ColaUntangle classified them into separate concerns, diverging from the dataset’s single-concern annotation. File formatting (3 cases) involve changes unrelated to human edits, such as automatic formatting. As shown in Figure 7 D, the first change in this commit is related to variable revise (not showed in Figure 7 D), the second change adds a newline at the end of the file, which is required by POSIX standards but inconsistently enforced by code editors (posixlinedef, ). Although such changes have no functional impact and are often ignored by developers, ColaUntangle classified them as a separate formatting concern, resulting in misclassification.
Since the datasets are built by selecting atomic commits and artificially tangling them, some standard atomic commits may exhibit high similarity in content or logic. This can cause ColaUntangle to misclassify them at a coarse grouping granularity. Specifically, we observed two common sub-types: similar change (11 cases) and mergeable logic (5 cases). Similar change refers to cases where two atomic commits contain highly similar modifications. As shown in Figure 8A, both commits modify different files with nearly identical changes to the return value of GetHashCode, leading ColaUntangle to group them as one concern. Mergeable logic occurs when changes in two atomic commits are tightly related. As shown in Figure 8B, one commit removes a registration statement, and the other re-adds it within an if block at the same location (line 47). ColaUntangle interpreted them as a single concern introducing conditional registration.
Overall, our approach occasionally untangles commits at an unexpected granularity, revealing challenges in both concern boundary modeling and dataset construction. In future work, we plan to improve concern modeling with human-in-the-loop interaction and develop more reliable methods for building commit untangling datasets to support this line of research.
7. Related Work
Early research on commit untangling primarily relied on heuristic rules. Barnett et al. (barnett2015Heuristic-rule, ) introduced the concept of diff-regions and proposed CLUSTERCHANGES to address the lack of automated support. Building on this, Herzig et al. (herzig2016impact, ) extracted relationships between code changes using the Confidence Voter and applied clustering algorithms to group diff-hunks. However, these approaches predominantly depended on static analysis techniques, limiting their applicability to dynamically typed languages.
Subsequent efforts introduced feature-based classification methods. Dias et al. (dias2015EpiceaUntangler, ) developed EpiceaUntangler, which mines relationships among code changes and employs a random forest classifier. This approach was suitable for untyped language environments where static analysis is inherently constrained. Later, ChangeBeadsThreader (yamashita2020changebeadsthreader, ), a streamlined variant of EpiceaUntangler, enabled developers to interactively refine clustering results.
More recent studies have shifted toward graph clustering combined with machine learning. Flexeme (parțachi2020flexeme, ) introduced Multi-version Program Dependency Graph ($\delta$-PDG) and Multi-version Name Flow Graph ($\delta$-NFG) and applied hierarchical clustering algorithms to untangle commits. SmartCommit (shen2021smartcommit, ) presented an interactive graph partitioning approach utilizing a scalable Diff Hunk Graph representation. It employed an edge shrinking algorithm to partition the graph into subgraphs and enabled developers to adjust outcomes.
Graph neural networks (GNNs) (li2015gated, ; xu2018powerful, ; kipf2016semi, ) have also been applied to untangling tasks. UTango (li2022utango, ) was the first method to incorporate GNNs, reusing Flexeme’s $\delta$-PDG while enriching it with node context information. By integrating clone detection, UTango overcame Flexeme’s limitations in handling implicit dependencies among cloned code. HD-GNN (fan2024HD-GNN, ) further advanced this line of work by proposing a fine-grained hierarchical graph model that captures the global context of code changes at both the entity and statement levels, addressing the problem of neglected hidden dependencies in prior approaches. More recently, Xu et al. (xu2025COMUNT, ) proposed an attributed graph modeling approach for detecting and untangling composite commits. Their method constructed a graph of code statements and dependencies, leveraging graph neural networks and clustering algorithms to achieve untangling.
8. Threats to Validity
Internal validity. One potential threat to internal validity arises from our definitions of explicit and implicit dependencies, which may not encompass all relationships between code changes that are crucial for untangling commits. To mitigate this, we leverage the semantic understanding, reasoning, and explanation capabilities of LLMs to infer dependencies beyond manually summarized rules. Furthermore, we provide worker agents with structured code change information, including both explicit and implicit contexts derived from $\delta$-PDG, to enable comprehensive reasoning over diverse dependency patterns. Another threat arises from the reliance of $\delta$-PDG on static analysis tools to model code. The accuracy and reliability of these tools can be limited, especially when handling complex language features, potentially leading to inaccurate analysis results.
External validity. A threat to external validity is potential data leakage, as the LLM may have been pre-trained on some open-source repositories used in our evaluation. However, this risk is minimal. The datasets are constructed by merging real atomic commits into new tangled commits, which are unlikely to appear in pretraining data. Moreover, the standalone LLM achieves only 60% and 63% $Accuracy^{c}$ on the C# and Java datasets, while our method improves this to 72% and 76%, suggesting the gain comes from our approach. Additionally, we attempted to prompt the LLM to reproduce exact commit histories but failed, indicating that the model does not memorize verbatim and relies on reasoning rather than recall. Another threat stems from the evaluation datasets themselves, as the tangled commits are constructed artificially by combining atomic commits. Although such datasets may include imprecise or non-representative data points, they are widely adopted in prior studies (parțachi2020flexeme, ; li2022utango, ; fan2024HD-GNN, ; xu2025COMUNT, ). In the future, we plan to develop more realistic and precise datasets to better evaluate commit untangling approaches.
9. Conclusion
In this paper, we propose ColaUntangle, the first LLM-driven collaborative model for commit untangling that jointly considers explicit and implicit dependencies among code changes. By integrating structured code change information with specialized agents in an iterative consultation process, ColaUntangle enables deeper semantic reasoning and interpretable decision-making. ColaUntangle achieves substantial performance gains, improving $Accuracy^{c}$ by 44% on the C# dataset and 100% on the Java dataset over the best-performing baseline. Our work introduces a new paradigm that simulates human-like collaborative reasoning, advancing the accuracy and transparency of automated commit untangling.