### v2

# LLM-Driven Collaborative Model for Untangling Commits via Explicit and Implicit Dependency Reasoning

Abstract.
Atomic commits, which address a single development concern, are a best practice in software development. In practice, however, developers often produce tangled commits that mix unrelated changes, complicating code review and maintenance. Prior untangling approaches—rule-based, feature-based, or graph-based—have made progress but typically rely on shallow signals and struggle to distinguish explicit dependencies (e.g., control/data flow) from implicit ones (e.g., semantic or conceptual relationships). In this paper, we propose ColaUntangle, a new collaborative consultation framework for commit untangling that models both explicit and implicit dependencies among code changes. ColaUntangle integrates Large Language Model (LLM)-driven agents in a multi-agent architecture: one agent specializes in explicit dependencies, another in implicit ones, and a reviewer agent synthesizes their perspectives through iterative consultation. To capture structural and contextual information, we construct Explicit and Implicit Contexts, enabling agents to reason over code relationships with both symbolic and semantic depth. We evaluate ColaUntangle on two widely-used datasets (1,612 C# and 14k Java tangled commits). Experimental results show that ColaUntangle outperforms the best-performing baseline, achieving an improvement of 44% on the C# dataset and 82% on the Java dataset. These findings highlight the potential of LLM-based collaborative frameworks for advancing automated commit untangling tasks.
Commit untangling, Explicit and implicit dependencies, LLMs, Multi-agent collaboration
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Software and its engineering Automatic programmingccs: Software and its engineering Software configuration management and version control systems
1. Introduction
In collaborative software development, a commit is the basic unit of code change, consisting of source file modifications and a message summarizing their purpose (shen2021smartcommit; dong2022fira; gomez2015visually). Ideally, each commit should be atomic, addressing a single concern—such as implementing a feature, fixing a bug, or refactoring (li2022utango; guo2019decomposing; ram2018makes; robillard2002concern). Atomic commits are a widely endorsed best practice for improving readability, facilitating reviews, and helping maintenance by clearly linking changes to their intention (tao2012empiricalwork; rigby2012contemporary; wang2019cora; herbold2022fine; baysal2016investigating). As a result, many software communities and companies recommend small, single-purpose commits (fan2024HD-GNN; linuxSubmittingPatches; microsoftGitGuidance).
However, in practice, developers often create tangled commits, which combine code changes related to multiple development concerns within a single commit (fan2024HD-GNN; murphy2011we). Tangled commits commonly arise due to time constraints or unclear boundaries between concerns during implementation (fan2024HD-GNN; parțachi2020flexeme; li2022utango). For example, a developer may simultaneously refactor code while fixing a bug alongside documentation updates, resulting in a commit that does not adhere to the principle of separation of concerns (herzig2016impact; herzig2013impact; nguyen2013filtering).
Empirical studies (herzig2013impact; herzig2016impact; tao2015partitioning; nguyen2013filtering) find that tangled commits are prevalent, indicating that 11% to 40% of commits in software repositories are tangled. Such commits hinder program comprehension, complicate maintenance (tao2012empiricalwork; parțachi2020flexeme; li2015gated). Moreover, tangled commits reduce the accuracy of automated tools such as bug prediction and localization models that rely on mining repository history (herzig2013impact; herzig2016impact; zhou2019devign; kirinuki2014hey; zimmermann2005mining; guo2017interactively). These tools often assume that each commit addresses a single concern, so tangled commits introduce noise and degrade model performance.
Researchers have proposed various approaches to untangling commits to split a tangled commit into separate coherent changes. These methods fall into three categories based on their automation level: heuristic rule-based, feature-based, and graph clustering-based. Heuristic methods use manually defined rules—such as line distance, textual similarity, or co-change frequency—to assess change relatedness (barnett2015Heuristic-rule; herzig2013impact; kirinuki2016splitting; wang2019cora). Feature-based methods employ handcrafted features (e.g., same method, class, or package) in supervised models to classify change pairs (dias2015EpiceaUntangler; yamashita2020changebeadsthreader). Graph clustering-based methods build program graphs, e.g., program dependency graphs (PDGs), to model structural dependencies and use clustering or node embeddings to group related changes (parțachi2020flexeme; li2022utango; xu2025COMUNT; fan2024HD-GNN).
Despite their methodological differences, these approaches share several key limitations. First, they often rely on surface-level signals or structural proximity and lack the capacity for deeper semantic reasoning. Graph-based methods attempt to encode some semantic relationships via graph embeddings, but they typically require substantial training and are limited in their ability to generalize across projects or languages. Second, most approaches act as black-box models, offering limited interpretability and little insight into why specific code changes are grouped together. Third and most critically, prior work does not clearly distinguish or integrate different types of dependencies among code changes. For instance, two edits may be connected via control flow (explicit) or may reflect a coherent conceptual refactoring (implicit), yet existing models either fail to recognize such links or treat them uniformly, resulting in reduced effectiveness on complex or ambiguous commits.
We argue that effective commit untangling must go beyond surface-level patterns and rely on a principled framework that explicitly accounts for both explicit and implicit dependencies. Explicit dependencies refer to observable relationships such as control or data flow, containment, or static code references. Implicit dependencies involve semantic connections—such as conceptual similarity, logical association, or shared intent—that are not necessarily reflected in the program structure. While some existing methods may partially capture these relationships through learned embeddings or co-change patterns, they do not explicitly model or directly leverage these dependencies as integral elements of the untangling process.
Recent advances in large language models (LLMs) have demonstrated strong performance in software engineering tasks that require semantic understanding, reasoning, and explanation (yang2024harnessing; fan2023large; hou2024large; fan2025exploring; li2024understanding). LLMs are particularly well-suited to identify implicit dependencies that go beyond what is detectable via static analysis or statistical patterns. However, a single LLM acting alone may not consistently balance different dependency perspectives or resolve ambiguous cases. Inspired by recent work on multi-agent collaboration (he2025llm; wang2025colacare; tang2023medagents; Kim2024mdagents), we propose decomposing the untangling task into subtasks, each handled by specialized agents, and enabling them to collaborate through consultation.
To this end, we introduce ColaUntangle, a collaborative consultation framework for commit untangling that integrates both explicit and implicit dependencies. We construct Explicit and Implicit Contexts to capture explicit and implicit dependencies between code changes. Then, we design a multi-agent architecture driven by LLMs: an explicit worker agent focused on explicit dependencies (e.g., control/data flow), an implicit worker agent focused on implicit dependencies (e.g., conceptual relationships), and a reviewer agent that synthesizes and reconciles their results. All agents provide not only untangling decisions but also explanations. Through iterative interaction, these agents simulate human-like consultation and collectively decide on the final untangling outcome.
To evaluate the effectiveness and efficiency of the proposed approach, we conduct experiments on the widely-used C# and Java datasets (parțachi2020flexeme; li2022utango; fan2024HD-GNN) with 1,612 and 14k tangled commits. The evaluation results show that ColaUntangle achieves better untangling results (i.e., 44% and 82% enhancement of effectiveness for C# and Java, compared to the best-performing baseline). Overall, our work makes the following contributions:
  • ColaUntangle: the first collaborative consultation model leveraging LLM-driven agents for commit untangling. We propose ColaUntangle, the first LLM-driven collaborative consultation model that untangles commits by considering explicit and implicit dependencies among code changes.
  • Formal and interpretable evidence framework. We define a clear and generalizable set of Explicit and Implicit Dependencies between code changes. This principled evidence framework provides direct guidance for LLM reasoning, which is foundational for explainable untangling.
  • A unified LLM framework for explainable, human-aligned untangling. We introduce a unified framework leveraging LLM-driven reasoning to infer semantics, robustly balance explicit and implicit dependencies, identify cosmetic edits and generate contextual explanations. Unlike prior black-box untangling models, this collaborative and human-centric architecture achieves explainable and interpretable untangling results.
  • Extensive empirical evaluation. We evaluate ColaUntangle against the state-of-the-art approaches, and the results demonstrate its superior performance. Beyond quantitative results, we conduct a detailed examination of both correctly untangled and misclassified cases, revealing how explainable reasoning contributes to accurate decisions and exposing the boundaries of current LLM-based untangling.
2. Motivation
High-quality software development demands atomic commits, where each commit addresses a single, isolated concern. However, in practice, developers frequently introduce tangled commits containing changes related to multiple independent activities. While prior work has attempted to untangle these commits, they often fall short due to an inability to fully capture the diverse and often subtle relationships between code changes. To define the scope of the problem and motivate the design of ColaUntangle, we analyze real-world commit examples, which crystallize four critical challenges faced by existing untangling methods.
2.1. Motivating Examples and Core Challenges
The examples in Figure 1, demonstrate the inadequacy of existing studies and highlight the necessity for an advanced, semantically-aware approach.
Figure 1. Motivating Examples
Figure 1A (Example 1) illustrates the changes made in commit 1ad6dbc of the Elasticsearch project. This atom commit updates the index used in the testMultipleBatchesWithLookupJoin function from "idx" to "no_sort_field_idx". The commit includes three separate code changes: lines 3–4 create the new index, lines 6–9 query using the updated index, and lines 11–12 delete the index. These changes do not exhibit any obvious data or control dependencies, nor do they share structural similarities in their syntax. However, they are semantically related, as they collectively form a complete and independent development activity involving the creation, usage, and deletion of an index. This semantic coherence indicates that they belong to a single concern. Existing untangling approaches may fail to handle such commits because they primarily rely on explicit data dependencies or structural similarities between code changes (parțachi2020flexeme; li2022utango; fan2024HD-GNN), overlooking the semantic relationships that often define a developer’s intent.
Challenge 1: Semantic Disconnect. Precisely untangling commits requires inferring the underlying semantic purpose and grouping changes that lack explicit structural links but form a cohesive, complete development activity.
Figure 1B (Example 2) shows another atom commit. Lines 3–6 define an enum class representing ordering options, and lines 8–16 implement a sorting method that utilizes this enum (only the if block is shown in the figure, as the else block has a similar structure). The code changes in lines 3–6 and lines 8–16 exhibit logical dependencies respectively, as they belong to the same class or method implementation. Furthermore, there is a clear data dependency: line 11 directly uses the value defined in the enum class introduced in lines 3–6. These changes are categorized as belonging to a single concern based on their logical and data dependencies, which are strong structural indicators for grouping.
However, many concerns are defined by implicit signals, where such strong explicit links are absent. Figure 1C (Example 3.) shows a tangled commit containing two distinct concerns: ”remove LicensingCharset” and ”reformat code”. The first concern (at lines 3 and 6) involves changes to the charset used in the TrialLicenseUtils class (only a subset of these changes is shown, as other charset modifications exhibit similar structure). Each of these changes is identical, replacing LicensesCharset with StandardCharsets. The affected statements are cloned code segments used to support different variables (issueTo and feature), resulting in high textual and structural similarity which is a crucial implicit signal for grouping. The variety of relationships between code changes—from the strong, explicit data flow in Example 2 to the subtle, implicit similarity here—presents an integration dilemma. Existing approaches struggle to meaningfully balance these two dependency types. They rely primarily on explicit dependencies, treating implicit signals (like similarity) as secondary factors or only using them for post-hoc adjustments (e.g., UTango (li2022utango)). This inability to fluidly and contextually integrate the entire spectrum of evidence (explicit and implicit dependency) leads to inaccurate or fragile untangling results.
Challenge 2: Contextual Integration. A unified approach must effectively integrate and weigh diverse forms of evidence—ranging from strong explicit dependencies (data/control flow) to subtle implicit relationships (semantics/similarity)—to accurately determine concern boundaries.
The second concern in Figure 1C (Example 3) focuses on reformatting code at lines 9, and 13, spanning two files. These changes represent typical cosmetic edits, which include syntactic formatting, refactoring, or non-functional textual modifications. Such edits are common in daily software development and usually belong to a single concern. However, existing approaches fail to detect and classify these changes correctly, as their rules (barnett2015Heuristic-rule; herzig2016impact), features (dias2015EpiceaUntangler), or graph representations (parțachi2020flexeme; li2022utango; fan2024HD-GNN) do not model cosmetic edits. Due to this oversight, they often fail to correctly isolate cosmetic edits, which can scatter across files and hide within larger functional concerns.
Challenge 3: Specialized Classification. Existing methods lack the capability to accurately identify and isolate non-functional cosmetic changes, leading to incomplete untangling.
Except for this, existing methods act as a black box, failing to provide the contextual reasoning behind their grouping decisions (parțachi2020flexeme; li2022utango; fan2024HD-GNN). This lack of a clear, contextual rationale is a significant limitation in a task requiring complex contextual judgment. A robust solution should offer interpretability, explaining its reliance on specific evidence (e.g., ”grouped due to semantic completion” or ”separated due to clear data independence”).
Challenge 4: Rationale and Explainability. Existing untangling methods lack the capacity to provide a clear, contextual rationale for their grouping decisions.
2.2. Key Principles
Based on the core challenges identified in Section 2.1, we propose ColaUntangle with an architecture founded on three key principles. These principles are designed to overcome the limitations by integrating deep semantic reasoning with a robust, multi-perspective decision-making process.
2.2.1. Principle 1: LLM-driven Foundation Models
To address Challenge 1 (Semantic Disconnect) and Challenge 3 (Specialized Classification), we decide to leverage the powerful capabilities of LLMs rather than traditional machine learning or neural network models to untangle commits. LLMs exhibit superior performance in comprehending code semantics, performing reasoning, and adapting to diverse decision-making tasks (liu2024large; jin2024llms). By employing LLM-driven agents, ColaUntangle can effectively understand the intent and implicit dependencies underlying code changes beyond purely structural or syntactic similarities. This enables our approach to untangle commits with a deeper semantic understanding and accurately classify non-functional edits which is difficult to achieve based on heuristic rules, feature representations and graph models.
2.2.2. Principle 2: Explicit and Implicit Dependency
Based on the motivating examples, we find that code changes can be grouped into different concerns due to various factors. To effectively guide the LLM’s reasoning, we formalize the diverse observations from our motivating examples into a structured framework of evidence. We categorize the characteristics linking code changes into two distinct, orthogonal types: Explicit Dependencies and Implicit Dependencies.
Explicit dependencies capture the direct relationships between code changes, such as data or control dependencies. In contrast, implicit dependencies capture the underlying reasons why code changes without obvious structural or data relationships may still belong to the same concern, such as semantic similarity or shared development purpose. Specifically, we define the following rules:
Explicit Dependencies Rules:
  • Code changes with data dependencies often belong to the same concern.
  • Code changes with control dependencies often belong to the same concern.
Implicit Dependencies Rules:
  • Code changes with semantic similarity may belong to the same concern.
  • Code changes with high textual or structure similarity may belong to the same concern.
  • Code changes introduced for cosmetic edits, such as syntactic formatting, refactoring, or non-functional textual modifications, often belong to the same concern.
Unlike prior studies that offered vague or indirect definitions of explicit and implicit dependencies, this formal set of rules provides a direct, comprehensive, and interpretable basis for untangling, enabling the LLM agents to reason directly over these untangling rules. The LLM agents can directly reason over these explicit principles. This design is fundamental to achieving both robust evidence integration (Challenge 2) and generating the clear, contextual rationale required for explainability (Challenge 4).
2.2.3. Principle 3: Collaborative Consultation
Recognizing that untangling requires resolving conflicts between the diverse evidence defined in Principle 2, and to address Challenge 2 (Contextual Integration) and Challenge 4 (Rationale and Explainability), we implement a multi-agent collaborative consultation process comprising distinct roles and iterative negotiation. Specifically, we define two types of worker agents: one specializing in analyzing explicit dependencies between code changes and the other focusing on implicit dependencies. Additionally, we introduce a reviewer agent that synthesizes and integrates the outputs from both worker agents. These three agents engage in multiple rounds of collaborative consultation, producing the final untangling result upon reaching consensus.
3. Methodology
Figure 2. Overall Workflow of ColaUntangle
We propose ColaUntangle, a collaborative framework that leverages LLM-driven agents to untangle commits by identifying both explicit and implicit dependencies among code changes. As shown in Figure 2, ColaUntangle first extracts structured code change information, including commit diffs and contexts derived from source code. It then employs a multi-agent consultation process, where agents iteratively analyze dependencies and synthesize untangling decisions until reaching consensus.
3.1. Structured Code Change Information
To enrich the contextual information available for dependency analysis, we first extract explicit contexts and implicit contexts from the program dependency graphs (PDGs) from the source code. Our method centers on building a consolidated graph structure sensitive to code changes, followed by targeted reduction and compression.
A program dependency graph (PDG) (ferrante1987PDG) is a directed graph that consists of a set of nodes N and a set of edges E, where each node n $\in$ N represents a program statement or a conditional expression; and each edge e $\in$ E represents the data flow or control flow among the statements within a single version of the code (ferrante1987PDG; le2014patch). However, PDG can only describe the dependency relationships between statements of individual program.
To simultaneously reflect statement dependencies before and after program modification, we first considered the capabilities of the Multi-version Program Dependency Graph ($\delta$-PDG) (parțachi2020flexeme; kim2006program), which is conceptually defined as the disjoint union of all nodes and edges from the PDGs of versions $i$ and $j$. We utilized this structural concept to construct our Commit-Level Merged Graph ($G_{Merged}$). However, $G_{Merged}$ is often massive, presenting a critical challenge for context-aware analysis with LLMs due to input length constraints, leading to inefficient token utilization and potentially diffusing the model’s focus away from the core changes. Therefore, we further explore the specialized extraction of contexts from $G_{Merged}$ to derive representations optimized for LLM consumption, namely the Explicit Context ($G_{Explicit}$) and the Implicit Context ($G_{Implicit}$).
We first construct the Commit-Level Merged Graph ($G_{Merged}$), which serves as the consolidated foundation for our context extraction. As illustrated in Algorithm 1, the construction process begins by taking the PDGs before ($\mathcal{G}_{B}$) and after ($\mathcal{G}_{A}$) the code change for every affected file, along with the commit’s line-level difference ($\mathcal{D}$). Using $\mathcal{D}$, we mark nodes in $\mathcal{G}_{B}$ corresponding to deleted lines as Deleted, and nodes in $\mathcal{G}_{A}$ corresponding to added lines as Added.
Following this initial marking, all file-level marked PDGs are aligned and merged into the single $G_{Merged}$ structure, where common unchanged nodes are consolidated using a fuzzy matching strategy (Lines 1-8). This construction relies on a robust fuzzy node matching strategy to identify semantically equivalent nodes between the two versions. Specifically, a node from $\mathcal{G}_{B}$ is matched with one from $\mathcal{G}_{A}$ if their node type, their structural context similarity, and their label content similarity all exceed defined fuzziness cutoffs ($\tau$), which is set to 1.0 following the practice of Flexeme (parțachi2020flexeme). Nodes successfully matched are consolidated as Unchanged Nodes. The resulting $G_{Merged}$ is constructed by consolidating the matched nodes (Unchanged Nodes) and their edges, retaining the unmatched nodes from $\mathcal{G}_{B}$ (Deleted nodes) and $\mathcal{G}_{A}$ (Added nodes), along with all their respective edges. This approach yields $G_{Merged}$, which represents the disjoint union of nodes and edges across both versions, with common elements consolidated, thereby effectively capturing the complete dependency scope of the committed changes.
Algorithm 1 Explicit and Implicit Contexts Generation
1: $\mathcal{D}$: Commit diff list (lines added/deleted); $\mathcal{G}_{B},\mathcal{G}_{A}$: PDGs of modified files (Before, After); $\tau$: Node matching fuzziness cutoffs.
2: $G_{Explicit}$: Explicit Context.
3: $G_{Implicit}$: Implicit Context.
[…]
5:for each modified file $F$ do
6:  $\mathcal{D}_{F}\leftarrow\text{DIFF\_FOR\_FILE}(\mathcal{D},F)$
7:  $G_{B,F}^{\prime}\leftarrow\text{MARK\_PDG}(\mathcal{G}_{B}[F],\mathcal{D}_{F})$ // Mark deleted nodes according to commit diffs
8:  $G_{A,F}^{\prime}\leftarrow\text{MARK\_PDG}(\mathcal{G}_{A}[F],\mathcal{D}_{F})$ // Mark added nodes according to commit diffs
9:  $G_{\Delta,F}\leftarrow\text{MERGE\_MARKED}(G_{B,F}^{\prime},G_{A,F}^{\prime},\tau)$ // Generate Merged PDG for each file
10:  $\mathcal{L}_{files}\leftarrow\mathcal{L}_{files}\cup\{G_{\Delta,F}\}$
11:end for
[…]
17:for each pair $(S,T)$ in $N_{Changed}\times N_{Changed}$ do // Explicit Context Generation
18:  if $S\neq T$ and $\text{PATH\_EXISTS}(G_{Merged},S,T)$ then
19:   $\text{ADD\_EDGE}(G_{Explicit},S,T,\text{type=`shortcut'})$
20:  end if
21:end for
22:return $G_{Explicit},G_{Implicit}$
For most modification operations, the changed parts account for only a small proportion of the entire program. To reduce unnecessary information, we introduce the concept of explicit contexts. The Explicit Context (i.e., $G_{Explicit}$) is a highly compressed graph designed to capture the logical dependency flow solely among the modified statements, disregarding intervening unchanged code. To construct $G_{Explicit}$, we first initialize a subgraph containing only the changed nodes ($N_{\text{Changed}}$), as shown in Algorithm 1 (Line 10). We then apply dependency compression by adding shortcut edges: for every distinct pair of nodes $(S,T)\in N_{\text{Changed}}$ that have a dependency path in the complete $G_{Merged}$, a direct shortcut edge is added between $S$ and $T$ in $G_{Explicit}$ (Lines 14-18). Although $G_{Explicit}$ loses all surrounding context, it effectively preserves the critical data and control dependencies between the modified statements, offering a minimal, dependency-centric view of the change impact.
Conversely, the Implicit Context (i.e., $G_{Implicit}$ ) is designed to retain neighbor code information without applying compression. $G_{Implicit}$ is generated by extracting the subgraph that includes all changed nodes ($N_{\text{Changed}}$) and their direct one-hop neighbors in $G_{Merged}$ (Lines 11-12). $G_{Implicit}$ strictly preserves only the original PDG edges present among these nodes and contains no shortcut edges. By including the immediate neighbors, $G_{Implicit}$ offers a larger, more comprehensive view of the surrounding code and structural environment than $G_{Explicit}$, enabling LLMs to leverage richer semantic information for accurate analysis.
Figure 3. Code Change Example with Explicit and Implicit Contexts
Figure 3 illustrates a code change scenario involving two statement replacements (at lines 2 and 6). Figure 3 (a) shows the $G_{Merged}$ of this code snippet, where red nodes represent the statements before the change, green nodes represent the modified statements, and gray nodes denote the unchanged statements. Figure 3 (b) presents the explicit contexts ($G_{explicit}$), which are obtained by compressing the $G_{Merged}$ in Figure 3 (a). Specifically, we retain the changed nodes and their direct dependency edges; additionally, if two changed nodes lack a direct dependency but are connected via a dependency path (channel), we add a compressed edge between them. For example, nodes 2 and 6 have no direct dependency but are connected via a path through nodes 4 and 5, resulting in a compressed edge between them. This compressed edge is labeled with the variables involved in the data flow along the original dependency path. Figure 3 (c) shows the implicit contexts ($G_{implicit}$), which retain both the changed nodes and their immediate neighbors (within one-hop dependencies). In this example, node 9 is excluded because it does not have a direct dependency with any of the changed nodes.
3.2. Multi-Agent Collaborative Consultation
As shown in Figure 2, the multi-agent collaborative consultation involves three types of agents: two worker agents specializing in mining explicit and implicit dependencies between code changes, respectively, and a reviewer agent responsible for reviewing and synthesizing the untangling results. The process begins with the two worker agents generating an initial untangling of a tangled commit based on its structured code information. The reviewer agent then synthesizes these results, initiating the collaborative consultation. In each iteration, the worker agents provide feedback on the current synthesized result, which the reviewer agent uses to revise synthesized result accordingly. This iterative process continues until the agents reach a consensus or a maximum number of rounds is reached.
3.2.1. Agent Building
(a) Worker Agent
(b) Reviewer Agent
Figure 4. Agents in Multi-Agent Collaboration
ColaUntangle builds three types of agents. The first two are worker agents, each specializing in untangling commits based on either explicit or implicit dependencies between code changes, as illustrated in Figure 4(a). The third is a reviewer agent, which specializes in reviewing and synthesizing the results from the worker agents, as shown in Figure 4(b).
Worker Agent. Based on the explicit and implicit dependency rules summarized in Section 2.2, we define two types of worker agents, as illustrated in Figure 4(a). The worker agents specialize in untangling commits by analyzing explicit and implicit dependencies between code changes, respectively. We refer to the explicit worker agent as EA and the implicit worker agent as IA. We define the roles of these agents using the following role definition prompt: “You are an expert in untangling tangled commits, specializing in analyzing explicit (or implicit) dependencies between code changes and comprehending their underlying semantic relationships”. We also provide each agent with the corresponding explicit or implicit dependency rules (concluded in Section 2.2) to guide their analysis. Each worker agent has two main functions. The first function is “untangle”. EA untangles tangled commits based on the inputs of commit diffs ($\mathcal{D}$) and explicit contexts ($G_{Explicit}$), while IA performs untangling based on commit diffs ($\mathcal{D}$) and implicit contexts ($G_{Implicit}$). Both agents are required to generate explanations for their decisions to facilitate understanding of their underlying criteria. The second function is “validate”. In this function, each worker agent compares the synthesized result generated by the reviewer agent (described in the next section) with its own untangling result and provides its opinion on the synthesized result. This includes indicating agreement or disagreement and, if disagreeing, providing a revised untangling result along with an explanation.
Reviewer Agent. The untangling results produced by the two worker agents may conflict due to their different expertise and the presence of diverse explicit and implicit dependencies within code changes. To resolve these conflicts, we define a reviewer agent, referred to as RA, which specializes in reviewing and synthesizing the untangling results from the worker agents, as illustrated in Figure 4(b). We define the reviewer agent’s role using the following role definition prompt: “You are an expert reviewer specializing in reviewing and synthesizing untangling results from worker agents. You have deep expertise in analyzing both explicit and implicit dependencies between code changes”. To support its role, we provide RA with both explicit and implicit dependency rules to guide its synthesis process. RA has three main functions: synthesize, revise and stop. In the synthesize function, RA synthesizes the results from EA and IA based on the structured code change information ($\mathcal{D}$, $G_{Explicit}$, and $G_{Implicit}$). In the revise function, RA revises its synthesized result based on the new opinions generated by the two worker agents during validation. Finally, in the stop function, when the collaborative consultation process reaches its maximum iteration limit, RA produces the final untangling result. For each function, RA is required to provide clear explanations for its decisions.
3.2.2. Untangling Workflow
The workflow of ColaUntangle is illustrated in Figure 2. The collaboration begins with the worker agents (EA and IA) providing their initial untangling results. Subsequently, the reviewer agent (RA) synthesizes these results into a unified untangling outcome. This initiates an iterative collaborative consultation process, during which EA and IA express their opinions on the current synthesized result. RA then considers their feedback and revises the synthesized result accordingly. This iterative process continues until EA and IA reach a consensus on RA’s synthesized result or a predefined maximum number of rounds is reached. Specifically, the untangling workflow comprises three phases: generation of initial reviews, synthesis of preliminary untangling results, and collaborative consultation process.
In the first phase, generation of initial reviews, EA and IA are instructed to generate their preliminary untangling results $r_{EA}$ and $r_{IA}$ using their Untangle function, based on the commit diffs ($\mathcal{D}$) and the explicit ($G_{Explicit}$) or implicit ($G_{Implicit}$) contexts:
$r_{EA}=EA.Untangle(Prompt(\mathcal{D},G_{Explicit}))$
$r_{IA}=IA.Untangle(Prompt(\mathcal{D},G_{Implicit}))$
In the second phase, synthesis of preliminary untangling results, RA uses its Synthesize function to review and integrate the results from EA and IA into a synthesized result $r_{RA}$
$r_{RA}=RA.Synthesize(Prompt(r_{EA},r_{IA},\mathcal{D},G_{Explicit},G_{Implicit}))$
Algorithm 2 Collaborative Consultation
1:Experts EAEA, IAIA, RARA; initial synthesized result $r_{RA}$; maximum attempts $t$; prompt template PromptPrompt
2:Final result $r_{f}$
[…]
4:while $consensus=False$ and $round<t$ do // Iterative consultation
5:  $round\leftarrow round+1$
6:  $consensus\leftarrow True$
7:  $EA\_opinion\leftarrow EA.validate(Prompt(r_{RA}))$
8:  $IA\_opinion\leftarrow IA.validate(Prompt(r_{RA}))$
9:  if EA_opinion.agree=FalseEA\_opinion.agree=False or IA_opinion.agree=FalseIA\_opinion.agree=False then
10:   $consensus\leftarrow False$
11:   if $round=t$ then
12:     $r_{RA}\leftarrow RA.stop(EA\_opinion,IA\_opinion)$
13:   else
14:     $r_{RA}\leftarrow RA.revise(EA\_opinion,IA\_opinion)$
15:   end if
16:  end if
17:end while
[…]
19:return $r_{f}$
The final phase, collaborative consultation process, is iterative as shown in Algorithm 2. During each round $j$, the worker agents EA and IA evaluate the current synthesized result and express their agreement or disagreement. In cases of disagreement, they are required to provide detailed rationales. To be specific, in round $j$-th:
$EA\_opinion^{j}=EA.validate(Prompt(r_{RA}^{j-1}))$
$IA\_opinion^{j}=IA.validate(Prompt(r_{RA}^{j-1}))$
If both agents reach unanimous agreement, the iterative process terminates (next loop condition is not satisfied). Otherwise, if the maximum round limit has not been reached, RA uses its Revise function to consider the opinions from EA and IA, refining the current synthesized result to produce a new synthesized result accordingly:
$r_{RA}^{j}=RA.revise(Prompt(EA\_opinion^{j},IA\_opinion^{j}))$
If the maximum round limit is reached without achieving consensus, RA uses its Stop function to finalize the synthesized result based on the current results and the agents’ opinions:
$r_{RA}^{j}=RA.stop(Prompt(EA\_opinion^{j},IA\_opinion^{j}))$
In the next round, the worker agents once again express their opinions towards the new synthesized result and try to convince the RA. The RA considers their statements and further revises the result until all worker agents reach a consensus or the maximum number of rounds is reached.
4. Experimental Setups
4.1. Datasets
We use two datasets that have been widely used in previous work: 1) A C# dataset used in (chen2022attributedgraph; herzig2016impact; parțachi2020flexeme; li2022utango; fan2024HD-GNN), which contains 1,612 tangled commits (each of which has $\leq$ 3 concerns) across 9 GitHub projects. 2) A Java dataset used in (li2022utango; fan2024HD-GNN). The Java dataset contains over 14k tangled commits in 10 GitHub repositories, where the number of concerns in a commit ranges from 2-32. Both datasets were constructed by first selecting atomic commits based on criteria such as temporal proximity, namespace similarity, and file coupling, and then artificially tangling them via git cherry-picking to simulate realistic tangled commit scenarios.
4.2. Baselines
Commit Untangling-specfic Baselines. The Baselines used for C# dataset are:
  • Barnett et al. (barnett2015Heuristic-rule): A heuristic rule-based approach considering def-use, use-use, and same-enclosing-method relations among code changes.
  • Herzig et al. (herzig2016impact): A heuristic rule-based approach combining various Confidence Voters and build triangle partition matrix to untangle the commits.
  • Flexeme (parțachi2020flexeme): A graph clustering-based approach builds a $\delta$-NFG (proposed in Flexeme) of commits and then applies agglomerative clustering on this $\delta$-NFG to untangle the commits.
  • $\delta$-PDG+CV (parțachi2020flexeme): A variant of Flexeme by applying Herzig et al.’s confidence voters directly to $\delta$-PDG (proposed in Flexeme).
  • UTango (li2022utango): A Graph clustering-based approach builds a $\delta$-PDG from commits and then applies GNN and agglomerative clustering on this $\delta$-PDG to untangle the commits.
  • HD-GNN (fan2024HD-GNN): A Graph clustering-based approach builds an entity reference graph from commits and then applies GNN to aggregate the embeddings of nodes and edges in the graph and classify the relation between the code changes.
The baselines used for Java dataset are: Barnett et al. (barnett2015Heuristic-rule), Herzig et al. (herzig2016impact), UTango (li2022utango), and HD-GNN (fan2024HD-GNN) mentioned above, as all these approaches support analysis of Java program files. SmartCommit (shen2021smartcommit) is a graph clustering-based interactive method that builds a Diff Hunk Graph and applies a graph partitioning algorithm to untangle commits. We also include SmartCommit and its baselines, Base-1 (a rule-based approach that groups all changes into a single concern) and Base-2 (which groups changes by file), in our comparison.
LLM-based Baselines. We incorporate baselines utilizing single LLM for both C# and Java datasets. We employ two different prompt strategies: zero-shot and zero-shot CoT. Zero-shot refers to prompting LLM to perform a task without providing examples, zero-shot CoT directly incorporates the prompt Let’s think step by step after a question to facilitate inference (kojima2022zeroshot).
4.3. Metrics
We use two evaluation metrics: $Accuracy^{c}$ proposed by Utango (li2022utango) and $Accuracy^{a}$ proposed by Flexeme (parțachi2020flexeme). $Accuracy^{c}$ is defined as the percentage of the changed statements that are labeled with a correct cluster/concern in all the statements in a commit:
$Accuracy^{c}=\frac{\#changed\ stmts\ w.\ correct\ clusters}{\#all\ changed\ stmts\ in\ commit}$
$Accuracy^{a}$ is similar to $Accuracy^{c}$ except that it considers all statements:
$Accuracy^{a}=\frac{\#stmts\ w.\ correct\ clusters}{\#all\ stmts\ in\ commit}$
4.4. Implementation Details
We construct the PDGs of C# and Java dataset with the assistance of the Flexeme extractor [^1] https://github.com/PPPI/Flexeme and Progex extractor [^2] https://github.com/ghaffarian/progex respectively. We utilize the widely adopted and publicly available DeepSeek-V3 model (liu2024deepseek) from the Azure OpenAI service [^3] https://learn.microsoft.com/en-us/azure/ai-services/openai/ as the reasoning engine in our experiments, which is also used for the LLM-based baselines. All experiments are conducted under a zero-shot setting. All LLMs used in our experiments are accessed via their online APIs, all decoding parameters are kept at their default values. From an initial evaluation on 50 randomly selected tangled commits, we observed an average of 1.27 consultation rounds to reach consensus. Thus, we set the maximum number of rounds $t$ to 3, aligning with ColaCare (wang2025colacare) (a medical multi-agent collaborative consultation model), and balancing iterative refinement with computational efficiency. In terms of computational cost, our approach incurs approximately \$2.92 per 100 examples (around ¢2.92 per example), with an average untangling time of 71.63 seconds per example (Context Building: 9.79 s; Generation of Initial Results: 12.23 s; Synthesis of Preliminary Untangling Results: 14.68 s; Collaborative Consultation Process: 34.93 s) [^4] We acknowledge that our approach requires longer inference time compared to prior methods. However, it eliminates the need for manual heuristic rule design, feature engineering, and model training, making it relatively cost-effective given its significant improvement in untangling accuracy.. We conduct experiments on a Linux server running Ubuntu 22.04 system, equipped with an Intel® Core™ i7 processor 14700k, 32GB RAM.
5. Experimental Results
To evaluate ColaUntangle, we seek to answer the following research questions:
  • [RQ1. Overall Performance] How well does ColaUntangle perform in comparison with the state-of-the-art commit untangling approaches?
  • [RQ2. Ablation Study] What is the contribution of each proposed module to the performance?
  • [RQ3. Sensitivity to Different LLMs] How does ColaUntangle’s performance vary with different LLMs?
Table 1. Performance on the C# Dataset ($Accuracy^{c}$$Accuracy^{a}$%)
Method CL CM HF HU LE NA NJ HI RS Overall
Previous Untangling Approaches
Barnett et al. 14/20 */20 12/15 13/25 7/16 */9 7/13 10/14 10/13 7/13
Herzig et al. 28/58 */65 28/62 27/53 27/66 */63 28/64 26/57 31/70 28/66
$\delta$-PDG + CV 34/81 */90 36/86 30/69 35/78 */83 34/78 37/94 33/64 35/84
Flexeme 34/87 */70 33/77 33/70 35/80 */87 27/62 32/80 34/86 34/83
UTango 46/90 */* 46/88 44/89 46/89 */* 41/83 46/91 46/88 46/90
HD-GNN 52/91 */* 52/90 50/90 50/91 */* 45/86 54/93 47/90 50/90
ColaUntangle Variants
$LLM_{ZeroShot}$ 62/90 68/90 64/93 44/66 58/92 52/88 68/97 63/91 60/92 60/90
$LLM_{ZeroShotCoT}$ 63/90 68/90 64/93 45/66 59/92 54/88 68/97 65/92 60/92 61/90
ColaUntangleno_comments 74/94 74/92 80/96 65/76 70/94 58/90 66/97 68/93 60/93 70/93
ColaUntangle 75/94 72/90 80/96 65/77 70/94 65/91 75/98 74/93 64/93 72/93
  • * refers no available data point.
    ColaUntangleno_comments refers to running our method on the dataset excluding comments.
Table 2. Performance on the Java Dataset ($Accuracy^{c}$%)
Method SB ES RJ GU RE DU GH ZX DR EB Overall
Previous Untangling Approaches
Barnett et al. 21 27 24 25 19 18 24 20 18 21 17
Herzig et al. 25 32 27 27 25 20 27 26 29 24 23
SmartCommit 29 34 31 33 29 24 33 34 32 27 30
Base-1 14 19 18 21 22 15 22 21 17 13 17
Base-2 23 22 28 31 24 19 27 29 29 22 24
UTango 32 36 33 37 34 30 35 38 34 31 34
HD-GNN 37 40 38 42 36 32 40 42 39 36 38
ColaUntangle Variants
$LLM_{ZeroShot}$ 59 66 57 47 47 57 51 65 64 55 59
$LLM_{ZeroShotCoT}$ 61 69 60 66 52 60 48 63 64 56 63
ColaUntangleno_comments 68 66 69 67 60 65 49 71 73 62 67
ColaUntangle 66 72 67 70 53 69 54 72 73 63 69
  • Since previous studies do not report $Accuracy^{a}$% on the Java dataset, we follow their practice and report only $Accuracy^{c}$% to ensure a fair comparison.
    ColaUntangleno_comments refers to run ColaUntangle on the dataset excluding comments.
5.1. RQ1: Overall Performance
Table 1 compares metrics on the C# dataset (including nine GitHub projects: CL: Commandline, CM: CommonMark, HF: Hangfire, HU: Humanizer, LE: Lean, NA: Nancy, NJ: Newtonsoft.Json, NI: Ninject, RS: RestSharp). Compared to previous untangling approaches, ColaUntangle outperforms SOTA, improving Barnett et al. (barnett2015Heuristic-rule), Herzig et al. (herzig2016impact), $\delta$-PDG+CV (parțachi2020flexeme), Flexeme (parțachi2020flexeme), Utango (li2022utango) and HD-GNN (fan2024HD-GNN) by 928.57%, 157.14%, 105.71%, 111.76%, 56.52% and 44.00% for $Accuracy^{c}$; 615.38%, 40.91%, 10.71%, 12.05%, 3.33% and 3.33% for $Accuracy^{a}$. For LLM-Based baselines, ColaUntangle improves $LLM_{ZeroShot}$ and $LLM_{ZeroShotCoT}$ by 16.67% and 16.39% on $Accuracy^{c}$, and by 3.33% on $Accuracy^{a}$ for both. Compared to ColaUntangleno_comments, which excludes comment statements, ColaUntangle achieves a 2.86% higher $Accuracy^{c}$, while $Accuracy^{a}$ remains the same.
Table 2 shows the Java dataset comparison (including ten GitHub projects: SB: spring-boot, ES: elasticsearch, RJ: RxJava, GU: guava, RE: retrofit, DU: dubbo, GH: ghidra, ZX: zxing, DR: druid, EB: EventBus), where ColaUntangle improves Barnett et al. (barnett2015Heuristic-rule), Herzig et al. (herzig2016impact), Base-1, Base-2, SmartCommit (shen2021smartcommit) Utango (li2022utango) and HD-GNN (fan2024HD-GNN) by 305.88%, 200.00%, 305.88%, 187.50%, 130.00%, 102.94% and 81.58% for $Accuracy^{c}$. For LLM-based baselines, ColaUntangle improves over $LLM_{ZeroShot}$ and $LLM_{ZeroShotCoT}$ by 16.95% and 9.52% respectively, and improves over ColaUntangleno_comments by 2.99% for $Accuracy^{c}$.
The results show that ColaUntangle consistently outperforms all baseline models, with notable improvements in the $Accuracy^{c}$ metric. Importantly, unlike most existing methods, ColaUntangle can untangle comment statements, which are common cosmetic edits in real-world commits, making it more practical for software development. ColaUntangleno_comments also surpasses all prior approaches, indicating that while comments can slightly enhance performance, our framework remains effective even without them. To be note, the single LLM’s overall accuracy outperforms commit untangling specific baselines, indicating the effectiveness of LLM in capturing the semantic information of tangled commit and executing the untangling task.
Regarding collaborative consultation rounds, the average number of rounds is 1.39: 75.44% of cases converge within one round, 12.11% within two rounds, and 13.55% reach the maximum of three rounds. Notably, in 48.91% of evaluated commits, the explicit and implicit worker agents initially produced non-consensual results, underscoring the importance of collaborative consultation for resolving divergent untangling perspectives.
(a) C# Repositories
(b) Java Repositories
Figure 5. Average $Accuracy^{c}$ and Number of Changed Statements in Tangled Commits of Each Repository
We also investigate the factors contributing to accuracy differences across repositories. As shown in Figure 5, for both the C# and Java datasets, the purple bars represent the average $Accuracy^{c}$ in each repository, while the red bars indicate the average number of changed statements of tangled commits in each repository. We observe a general trend where $Accuracy^{c}$ decreases as the number of changed statements increases. This may be because abundant changes in the commit cause LLMs forgetting context or untangling commits more finely than humans expect. These findings suggest future work on hierarchical prompting or human-in-the-loop approaches to better align with human expectations.
Answer to RQ1: ColaUntangle demonstrates improvements over prior baselines on both datasets, improving $Accuracy^{c}$ by 44% on the C# dataset and 82% on the Java dataset over the best existing baseline. The results highlights its potential for practical application in untangling commits.
5.2. RQ2: Ablation Study
We conduct an ablation study to evaluate the contribution of each component within ColaUntangle. As shown in Table 3, we define four ablation settings: (1) w/o Info Tools: Removes explicit and implicit contexts extracted from the PDGs. (2) w/o Collaborative Consultation: Uses a single LLM model to mine both explicit and implicit dependencies without the multi-agent consultation process. (3) w/o Explicit Worker Agent (W1) and Collaborative Consultation: Retains only the implicit worker agent to perform untangling. Since there is only a single agent remaining, the collaborative consultation process becomes unnecessary, and the reviewer agent is thus excluded. (4) w/o Implicit Worker Agent (W2) and Collaborative Consultation: The converse of (3), retaining only the explicit worker agent.
From the results, we observe that excluding Collaborative Consultation (setting 2) leads to the most significant performance degradation (16.90%) [^5] For the sake of brevity, all mentions of performance degradation in this section refer specifically to a drop in $Accuracy^{c}$, as it better reflects the model’s performance in real-world tasks compared to $Accuracy^{a}$., highlighting its critical role in achieving optimal untangling results. Removing either the explicit or implicit worker agent (settings 3 and 4) also results in significant performance drops (18.99% and 19.72%, respectively). However, it is important to note that these two settings inherently exclude Collaborative Consultation due to the presence of only a single agent, making their performance degradation partly attributable to the absence of collaborative consultation. When we isolate the effect by excluding the performance drop caused by the lack of collaborative consultation, settings 3 and 4 show additional reductions of only 2.09% and 2.82%, respectively. This indicates that while incorporating both explicit and implicit worker agents enhances performance, the collaborative consultation mechanism itself contributes the most substantial improvement in our framework. Removing Info Tools (setting 1) results in a performance decrease of 8.45%, suggesting that the structured context improves our model’s untangling effectiveness. However, compared to the substantial improvements brought by collaborative consultation, Info Tools mainly serves as foundational support by providing essential contextual information rather than driving the core reasoning and decision-making process.
Answer to RQ2: The ablation results suggest that while structured context information and the integration of both explicit and implicit worker agents supports untangling accuracy, the collaborative consultation mechanism play a more critical role in achieving the effectiveness of ColaUntangle.
Table 3. Ablation Study Results for Each Module ($Accuracy^{c}$$Accuracy^{a}$%)
Method C# Dataset Java Dataset Overall
w/o Info Tools 64/91 65/91 65/91
w/o Collaborative 60/90 59/91 59/91
w/o W1 w/o Collaborative 57/90 54/90 56/90
w/o W2 w/o Collaborative 58/90 55/92 57/91
ColaUntangle 72/93 69/93 71/93
  • W1(2) refer to worker agent that expert in discovering explicit (implicit) dependency between code changes.
5.3. RQ3: Sensitivity to Different LLMs
To evaluate the generalization capability of our framework, we include both closed-source proprietary models (GPT-4o, GPT-4o-mini from OpenAI, Claude-4-sonnet from Anthropic) and open-source frontier models (DeepSeek-V3 and Qwen3-235b) in our comparison. We compare ColaUntangle instantiated with these five models respectively. The performance presented in Table 4 shows that all these LLMs can effectively untangle commits, with DeepSeek-V3 performing slightly better. This suggests that while ColaUntangle is generally robust to the underlying LLM choice, stronger reasoning models still yield marginal performance gains. Furthermore, we analyze the average number of collaborative rounds required for consensus: DeepSeek-V3 (1.39), GPT-4o-mini (1.65), GPT-4o (1.34), Qwen3-235b (1.41), and Claude-4-sonnet (1.73). Notably, DeepSeek-V3 not only attains leading accuracy but also requires fewer consensus rounds compared to most competing models (except GPT-4o), highlighting its efficiency in balancing performance and computational cost.
Answer to RQ3: Our evaluation demonstrates that while all tested LLMs effectively handle commit untangling, DeepSeek-V3 achieves the optimal balance of accuracy and efficiency (fewer consensus rounds).
Table 4. Performance of Different LLMs ($Accuracy^{c}$$Accuracy^{a}$%)
Method C# Dataset Java Dataset Overall
GPT-4o-mini 72/93 64/91 69/93
GPT-4o 68/92 67/93 67/93
Qwen3-235b 71/93 66/92 69/92
Claude-4-sonnet 70/93 68/92 69/93
ColaUntangle (DeepSeek-V3) 72/93 69/93 71/93
6. Discussion
Although ColaUntangle performs well overall, examining both misclassified and correctly untangled commits provides a deeper understanding of its strengths and limitations, informing potential directions for future work.
6.1. Error Cases
In the RQ1 experiments, although ColaUntangle performed well overall, it still misclassified some cases. To support future improvements, we conducted a manual analysis following Fan et al. (fan2024HD-GNN), randomly sampling 50 misclassified cases. As shown in Table 5, we grouped the errors into two main types: fine and coarse grouping granularity. Fine grouping granularity refers to ColaUntangle untangles a tangled commit into concerns at finer granularity than the standard answer. In contrast, coarse grouping granularity refers to ColaUntangle untangles a tangled commit into concerns at coarser granularity than the standard answer. Further analysis identified five sub-types that specify the scenarios of these errors.
Table 5. Type of Error Cases
Type Sub-Type Number
Fine Grouping Granularity Judgment Error 14
Composite Commit 17
File Formatting 3
Coarse Grouping Granularity Similar Change 11
Mergeable Logic 5
Figure 6. Examples of Error Cases Related to Fine Grouping Granularity
For fine grouping granularity, we observed three representative sub-types: judgment errors (14 cases), composite commit (17 cases), and file formatting (3 cases). Judgment errors (14 cases) occur when ColaUntangle misinterprets the relationship between changes. For example, in Figure 6 A, changes across two files belong to a single concern related to signature revision, but ColaUntangle misclassified them into two concerns by treating lines 9–10 as an isolated function call modification (configure) unrelated to the other file. Similarly, in Figure 6 B, changes at lines 3–5 and 8–9 belong to the same concern (add more logging to jgroups disco), but ColaUntangle separated them because it focused on different logger methods (debug vs. warn). Composite commit (17 cases) occur when the standard atom commit in the dataset actually contains multiple unrelated activities. For instance, Figure 6 C shows a commit that modifies comments for the create function (lines 3–5) and changes the signature of canAcquire, with the two code changes separated by over 240 lines and lacking any logical connection. ColaUntangle classified them into separate concerns, diverging from the dataset’s single-concern annotation. File formatting (3 cases) involve changes unrelated to human edits, such as automatic formatting. As shown in Figure 6 D, the first change in this commit is related to variable revise (not showed in Figure 6 D), the second change adds a newline at the end of the file, which is required by POSIX standards but inconsistently enforced by code editors (posixlinedef). Although such changes have no functional impact and are often ignored by developers, ColaUntangle classified them as a separate formatting concern, resulting in misclassification.
Figure 7. Examples of Error Cases Related to Coarse Grouping Granularity
Since the datasets are built by selecting atomic commits and artificially tangling them, some standard atomic commits may exhibit high similarity in content or logic. This can cause ColaUntangle to misclassify them at a coarse grouping granularity. Specifically, we observed two common sub-types: similar change (11 cases) and mergeable logic (5 cases). Similar change refers to cases where two atomic commits contain highly similar modifications. As shown in Figure 7A, both commits modify different files with nearly identical changes to the return value of GetHashCode, leading ColaUntangle to group them as one concern. Mergeable logic occurs when changes in two atomic commits are tightly related. As shown in Figure 7B, one commit removes a registration statement, and the other re-adds it within an if block at the same location (line 47). ColaUntangle interpreted them as a single concern introducing conditional registration.
Overall, our approach occasionally untangles commits at an unexpected granularity, revealing challenges in both concern boundary modeling and dataset construction. In future work, we plan to improve concern modeling with human-in-the-loop interaction and develop more reliable methods for building commit untangling datasets to support this line of research.
6.2. Correct Cases
To further validate the effectiveness of our approach, we randomly selected a representative case from the dataset and conducted a comparative evaluation of ColaUntangle, UTango (li2022utango), and a single LLM. We selected UTango as a key comparator because it partially shares common structural data sources (utilizing $\delta$-PDG, which informs our structured code change) and its methodology touches upon the concepts of explicit and implicit factors. Specifically, UTango employs a sequential, two-stage process: extracting feature embeddings from the $\delta$-PDG for initial clustering (explicit context), followed by a code-clone detection process to refine groupings based on similarity (implicit dependency). This sequential approach is a critical point of contrast with our collaborative mechanism. We selected the single LLM (DeepSeek-V3) as a second comparator to isolate and intuitively observe the performance of LLM on direct reasoning. This baseline operates without access to the structured explicit/implicit contexts or the rule-based guidance provided to our multi-agent framework. Through this comparative setting, we aim to highlight the contribution of ColaUntangle’s multi-agent reasoning and context-aware mechanisms to improving untangling accuracy and interpretability.
Figure 8. Untangling Result of Different Model
As shown in Figure 8, the commit contained three distinct changes: a file update, and the addition of two new functions, SetJobParameter and CreateJob. While UTango and the single LLM failed to untangle this commit correctly, our framework successfully generated the correct answer, which is not a trivial task due to the subtle dependencies between the code changes. UTango failed for two key reasons: it could not identify the first code change because the $\delta$-PDG representation was incomplete (lacking the import statement at the file’s beginning), and it incorrectly grouped the two new functions together based on a structural similarity detected by its code clone detection process, ignoring their functional independence. Similarly, the single LLM also grouped the two functions together, explaining its decision based on a generic principle of separating “cosmetic edits from functional changes,” but it failed to discern the distinct purposes of the two new functions.
Figure 9. Untangling Process of ColaUntangle
The effectiveness of ColaUntangle lies in its underlying process, as shown in our step-by-step analysis (Figure 9). The Explicit Agent, using the explicit contexts and original commit diff, correctly identified the two new functions as separate entities, while the Implicit Agent’s semantic analysis noted their structural and semantic similarities. The Reviewer Agent then synthesized and balanced these two viewpoints, ultimately generating a correct grouping that preserved the functional independence of the new functions. The collaborative consultation process then led all agents to agree on this correct grouping. This example illustrates the combined power of our key ideas: the LLM’s superior semantic understanding, its ability to deal with cosmetic edits and other non-functional changes, and its capacity to use the explicit and implicit contexts as an active component to infer subtle dependencies. The collaborative framework ensures a balanced and accurate result by reconciling the distinct explicit and implicit viewpoints, which together make ColaUntangle’s performance superior to that of prior work and single-model baselines.
7. Related Work
Early research on commit untangling primarily relied on heuristic rules. Barnett et al. (barnett2015Heuristic-rule) introduced the concept of diff-regions and proposed CLUSTERCHANGES to address the lack of automated support. Building on this, Herzig et al. (herzig2016impact) extracted relationships between code changes using the Confidence Voter and applied clustering algorithms to group diff-hunks. However, these approaches predominantly depended on static analysis techniques, limiting their applicability to dynamically typed languages.
Subsequent efforts introduced feature-based classification methods. Dias et al. (dias2015EpiceaUntangler) developed EpiceaUntangler, which mines relationships among code changes and employs a random forest classifier. This approach was suitable for untyped language environments where static analysis is inherently constrained. Later, ChangeBeadsThreader (yamashita2020changebeadsthreader), a streamlined variant of EpiceaUntangler, enabled developers to interactively refine clustering results.
More recent studies have shifted toward graph clustering combined with machine learning. Flexeme (parțachi2020flexeme) introduced Multi-version Program Dependency Graph ($\delta$-PDG) and Multi-version Name Flow Graph ($\delta$-NFG) and applied hierarchical clustering algorithms to untangle commits. SmartCommit (shen2021smartcommit) presented an interactive graph partitioning approach utilizing a scalable Diff Hunk Graph representation. It employed an edge shrinking algorithm to partition the graph into subgraphs and enabled developers to adjust outcomes.
Graph neural networks (GNNs) (li2015gated; xu2018powerful; kipf2016semi) have also been applied to untangling tasks. UTango (li2022utango) was the first method to incorporate GNNs, reusing Flexeme’s $\delta$-PDG while enriching it with node context information. By integrating clone detection, UTango overcame Flexeme’s limitations in handling implicit dependencies among cloned code. HD-GNN (fan2024HD-GNN) further advanced this line of work by proposing a fine-grained hierarchical graph model that captures the global context of code changes at both the entity and statement levels, addressing the problem of neglected hidden dependencies in prior approaches. More recently, Xu et al. (xu2025COMUNT) proposed an attributed graph modeling approach for detecting and untangling composite commits. Their method constructed a graph of code statements and dependencies, leveraging graph neural networks and clustering algorithms to achieve untangling.
It is worth noting that GNN-based approaches, particularly UTango, have achieved promising results but still face certain intrinsic constraints that ColaUntangle aims to address. Specifically, UTango relies on GNNs operating over the $\delta$-PDG, where feature extraction is primarily based on structural proximity. While this design effectively captures local structural patterns, it limits the model’s ability to incorporate non-structural semantic context that is essential for fully understanding developer intent. As a result, UTango may face challenges in capturing semantic cohesion (Challenge 1) and in distinguishing specialized non-functional changes such as cosmetic edits (Challenge 3). Moreover, UTango integrates evidence in a sequential manner—first through embeddings of the $\delta$-PDG for grouping and then through clone detection for refinement—making it difficult to achieve the concurrent and context-aware balancing required for complex concerns (Challenge 2). Finally, as a black-box neural model, UTango provides limited interpretability regarding the rationale behind its grouping decisions (Challenge 4). In contrast, ColaUntangle leverages a multi-agent LLM architecture that facilitates deeper semantic inference, more integrated evidence fusion, and native generation of contextual explanations.
8. Threats to Validity
Internal validity. One potential threat to internal validity arises from our definitions of explicit and implicit dependencies. Our framework categorizes relationships as either explicit or implicit, and while explicit rules are obvious (data and control dependency), it’s impossible to manually enumerate every complex, context-dependent implicit semantic relationship that drives developer intent. To mitigate this, we shift our approach from a brittle, rule-based system to a flexible, reasoning-based paradigm. We acknowledge that manually summarized rules cannot cover all implicit cases. Instead, we perform extensive analysis of literature and code to define patterns covering the majority of critical implicit dependencies. We then provide our LLM-driven worker agents with both structured rules and rich contextual information derived from PDGs. Leveraging the LLM’s semantic understanding and reasoning capabilities, the implicit worker agent is able to infer latent dependencies that extend beyond these manually formalized rules. The subsequent collaborative consultation mechanism, where the reviewer agent reconciles the distinct views of the explicit and implicit workers, ensures a comprehensive analysis over diverse dependency patterns. The strong performance of our method suggests that our implicit dependency coverage, combined with the LLM’s inference power, is sufficient for robust practical performance. Another threat arises from the reliance of PDGs on static analysis tools to model code. The accuracy and reliability of these tools can be limited, especially when handling complex language features, potentially leading to inaccurate analysis results.
External validity. A threat to external validity is potential data leakage, as the LLM may have been pre-trained on some open-source repositories used in our evaluation. However, this risk is minimal. The datasets are constructed by merging real atomic commits into new tangled commits, which are unlikely to appear in pretraining data. Moreover, the standalone LLM achieves only 60% and 59% $Accuracy^{c}$ on the C# and Java datasets, while our method improves this to 72% and 69%, suggesting the gain comes from our approach. Additionally, we attempted to prompt the LLM to reproduce exact commit histories but failed, indicating that the model does not memorize verbatim and relies on reasoning rather than recall. Another threat stems from the evaluation datasets themselves, as the tangled commits are constructed artificially by combining atomic commits. Although such datasets may include imprecise or non-representative data points, they are widely adopted in prior studies (parțachi2020flexeme; li2022utango; fan2024HD-GNN; xu2025COMUNT). In the future, we plan to develop more realistic and precise datasets to better evaluate commit untangling approaches.
9. Conclusion
In this paper, we propose ColaUntangle, the first LLM-driven collaborative model for commit untangling that jointly considers explicit and implicit dependencies among code changes. By integrating structured code change information with specialized agents in an iterative consultation process, ColaUntangle enables deeper semantic reasoning and interpretable decision-making. ColaUntangle achieves substantial performance gains, improving $Accuracy^{c}$ by 44% on the C# dataset and 82% on the Java dataset over the best-performing baseline. Our work introduces a new paradigm that simulates human-like collaborative reasoning, advancing the accuracy and transparency of automated commit untangling. Our tool and data are available at https://github.com/ColaUntangle/ColaUntangle.